### v2

# Emergence in graphs with near-extreme constraints

Abstract.
We consider optimal graphons associated with extreme and near-extreme constraints on the densities of edges and triangles. We prove that the optimizers for near-extreme constraints are unique and multipodal and are perturbations of the previously known unique optimzers for extreme constraints. This proves the existence of infinitely many phases. We determine the podal structures in these phases and prove the existence of phase transitions between them.
1. Results and setting
1.1. New results
We investigate the large scale structures called podes which emerge as node number increases in simple graphs with given edge and triangle densities $(e,t)$. In particular we focus on phase transitions which appear as discontinuities in the pode structure as $(e,t)$ is varied.
The boundary curves in Figure 1 show the extreme accessible values of pairs $(e,t)$ of those densities. Using graphons, the Large Deviation Principle (LDP) of $\mathbb{G}(n,p)$ graphs, and a Boltzmann entropy taken from statistical mechanics [37], one can associate [37] with each extreme-density pair $(e,t)$ a unique optimal graphon, i.e. a graphon with the given densities which optimizes the LDP rate function.
Figure 1. The Razborov triangle, made from curves displaying the extreme values of pairs of accessible edge and triangle densities. The curvature of the “scallops” on the lower right is exaggerated for visibility.
In this paper we prove that the unique optimal graphons at the extreme densities extend smoothly into unique optimal graphons in the nearby interior of the triangle, and we fully determine their pode structure. An important technical result of the paper is a new technique for proving the existence of these near-extreme phases, in particular distinct phases emanating from each of the infinitely many scallops at the bottom of Figure 1.
1.2. The background
To prove sharp transitions in large simple graphs we depend heavily on the graph-limit or graphon formalism of Lovász et al [7, 8, 26, 27, 28] (following the work of Aldous [1], Hoover [18] and Kellenberg [20]; see [4, 16] for background). Graphons are an infinite-node idealization of graphs. They do not have edges, but they do have well-defined densities of finite graphs, including edges. Emergence contrasts a large scale with a small scale in finite simple graphs. The smallest scale is that of individual edges, and the large scale is that of dense graphs with many edges on many nodes. Our proofs use graphons in place of large finite graphs, but we justify the distinction. We prove the existence of arbitrarily sharp transitions in asymptotically large finite graphs, based on a dependency between edges produced by triangles. In 2010 Chatterjee and Varadhan proved [13] their LDP for $\mathbb{G}(n,p)$, built on top of the graphon formalism. We introduced in [37] the Boltzmann entropy $\mathbb{B}(e,t)$ as the exponential rate of growth of the number of graphs with densities $(e,t)$. Using the rate function $I(g)$ of the LDP, we proved [37] the representation of $\mathbb{B}(e,t)$ as the maximum of $S(g)=-I(g)$ over all graphons $g$ with edge-triangle constraints $(\varepsilon(g),\tau(g))=(e,t)$. The LDP then implied that if this constrained optimization has a unique graphon solution $g_{e,t}$ (an important proviso!), then all but an exponentially small fraction of large but finite graphs with edge and triangle densities close to $(e,t)$ are well-approximated by $g_{e,t}$ [38]. The latter implies that any true singularites proven at the infinite-node graphon scale represent the approximate phenomenon for exponentially-most large finite constrained graphs.
Completion of Figure 1 in 2012 [35] was a significant achievement by Pikhurko and Razborov. This paper is part of a series based on analyzing phase transitions which emerge at various places in the Razborov triangle.
Figure 2. Schematic drawing of a conjecture from 2017 [21], based on computer simulations of optimal graphons associated with the phases of large graphs with edge and triangle constraints.
In the initial papers we developed methods to determine a few phases far in the interior of the Razborov triangle, near the curve $t=e^{3}$ associated with graphs with independent edges. We defined a phase as a connected open set in the $(e,t)$ plane in which the unique optimal graphons, and therefore $\mathbb{B}$, vary smoothly with $(e,t)$, so that phase transitions would be seen as singularities separating phases. This assumed that smoothness prevailed over most of Figure 1 so one could concentrate on the transitions. Although this turned out to be the case, the proof took several years.
Our project thus amounted to showing that, apart from rare (codimension-1) values of $(e,t)$, the $S(g)$-maximizing graphon with densities $(e,t)$ is unique and varies smoothly with $(e,t)$, implying that $\mathbb{B}(e,t)$ is smooth at $(e,t)$. From computer simulations we were led in 2017 to conjecture [21] that this is the case except across the curves shown in Figure 2. Consequently, all unconstrained subgraph densities associated with exponentially-most graphs should be piecewise-smooth functions of $(e,t)$, with singularities at phase boundaries. This conjecture gave the project a more concrete goal.
Progress was slow until now. We first proved that the picture in Figure 2 was correct just above the Erdős-Rényi curve, $t=e^{3}$, (the $F(1,1)$ phase in the figure) [22], then just below the Erdős-Rényi curve when $e>1/2$ (the $B(1,1)$ phase) [32], and then a neighborhood of the $e=1/2$ line for $t<e^{3}$ (part of the $A(2,0)$ phase) [39, 33].
1.3. Connections with other optimization problems
By far the most highly developed mathematical formalism concerned with emergence is equilibrium statistical mechanics. In our project we have sought to create an analogous formalism for simple constrained graphs, motivated by the Razborov triangle, Figure 1. One of our goals is to better understand constrained graphs, to extend the classic study of extremal graphs. Another is to provide an example for other optimization problems facing similar obstacles. For this reason we give a brief sketch of emergence in statistical mechanics [41, 47, 46], emphasizing the significance of convexity.
A common feature of edge-triangle graphs and statistical mechanics is constrained optimization. There are several ways to view equilibrium statistical mechanics as a constrained optimization problem [19, 17, 40, 23], all of which involve the global optimization of any of a range of free energy functionals, or the entropy. The entropy in statistical mechanics is a measure of the number of possible particle configurations with given constraints. It is a fundamental quantity. It is no exageration to view statistical mechanics as built on the convexity of this entropy (see the lectures of Lanford in [23] and the introduction by Wightman in [19]). The convexity of the entropy allows one to analyze the system without loss of information by the use of a variety of free energies [46], such as the Gibbs free energy $G(p,T)$, which is more familiar than the entropy.
Figure 3. This is a crude sketch of the phases of bulk matter, separated by transition curves. There are more than 20 known different solid phases of water, different crystalline structures.
Figure 3 shows a primitive thermodynamic phase diagram, illustrating the pattern of solid and fluid phases in a physical bulk material (i.e. the large scale in the emergent picture), as functions of constraint parameters pressure $p$ and temperature $T$ [36]. The experimentally measurable Gibbs free energy $G(p,T)$ [47, 42] is found to vary smoothly within each region, but shows singular behavior when constraints cross some lower dimension curves (see [40] and section VI in [19]), where bulk material properties such as mass density and heat capacity can change abruptly.
In the analogy between constrained graphs and statistical mechanics, discussed for instance in [34], [44] and [12], edges play the role of individual particles and the energy of a graph is the number of copies of some subgraph $H$, such as triangles. A key to understanding phases in both statistical mechanics and constrained graphs is defining an order parameter [3], which is a function that is identically zero in one phase but nonzero in another. We show its use in Section 4.6.
Unfortunately, with graphs there is no analog of equivalent free energies. This is analyzed in Section 6. This was the starting point of our project in 2013 [37]. (The “inequivalence of ensembles” in constrained graphs was rediscovered and analyzed later; see [15].) This inequivalence presents a serious obstacle. In statistical mechanics free energies provide considerable technical advantages. In their absence we had to develop replacement tools. One such tool, which we introduce in this paper, is the “worth” $W(C)$ that we associate to each column $C$ of a graphon. Developing our variational arguments at the level of columns of graphons is a significant advance in our understanding of optimal graphons and therefore phase transitions in constrained graphs.
We note that some important optimization problems suffer from nonconvexity complications. For instance, the form of optimal transport theory developed by Kantorovich and Brenier was based on convex analysis. It required significant developments over several years to allow the original nonconvex Monge problem to access that convex analysis. (For an introduction to optimal transport see the preface in [45] or chapter 1 in [29]).
The emergence of structure in large but finite physical materials is dramatic, particularly the diversity of solid phases, such as the graphite and diamond phases of carbon. All the richness displayed by the phases of all materials, not just the pure elements but also the huge number of compounds like water and alcohol, is created by the electromagnetic interaction within and between molecules. All of inorganic chemistry comes down to the 100 or so different (integer) electric charges of atomic nuclei.
Emergence gives rise to the complicated structure of “water” in Figure 3 and in the edge-triangle model in Figure 2. The study of the emergence of such diversity from the interaction of invisibly small components of a small number of types has led to a great deal of interesting mathematics in the past twenty years. It is this promising history which was the motivation to bring the richness of statistical mechanics, which is built on the convexity of its entropy, into the nonconvex setting of the Boltzmann entropy $\mathbb{B}$ of constrained graphs.
1.4. Definitions
In order to state our results more precisely, we need a series of definitions. For an in-depth presentation of graphons see [25]. For an introduction to Large Deviation Principles (and also a compact introduction to graphons) see [9].
Here are a few of the basic definitions. A graphon is a measurable function $g:[0,1]^{2}\to[0,1]$, with $g(y,x)=g(x,y)$ for every ordered pair $(x,y)$. Two graphons $g_{1}$ and $g_{2}$ are said to be equivalent if there is a measure-preserving transformation $\sigma:[0,1]\to[0,1]$ such that $g_{2}(x,y)=g_{1}(\sigma(x),\sigma(y))$ for almost every $(x,y)$. The equivalence classes of graphons are called reduced graphons. The Shannon entropy, edge density and triangle density of a graphon are given by the functionals
[…]
where
(4) $H(u)=-[u\ln(u)+(1-u)\ln(1-u)]$
is the usual entropy of independent coin flips with probability $u$ of getting heads. Note that if $g_{1}$ and $g_{2}$ are equivalent graphons, then $S(g_{1})=S(g_{2})$, $\varepsilon(g_{1})=\varepsilon(g_{2})$ and $\tau(g_{1})=\tau(g_{2})$. These quantities should really be viewed as functionals on the space of reduced graphons. $S(g)$ is the negative of the LDP rate function $I(g)$.
The following definitions are less standard.
If the graphon $g$ maximizes $S(g)$ subject to constraints on the densities on edges and another graph, we say that $g$ is an optimal graphon at those densities. When we speak of an optimal graphon being unique, we mean that the reduced graphon is unique. In other words, that the optimal graphon is unique up to equivalence.
A graphon is said to be k-podal if we can partition the interval $[0,1]$ into measurable sets $I_{1},\ldots,I_{k}$ such that $g(x,y)$ is constant on each “rectangle” $I_{i}\times I_{j}$. By picking an appropriate representative of the equivalence class of $g$, we can assume that each $I_{i}$ is an interval. We refer to the sets $I_{i}$ as podes. If $g$ is $k$-podal for some integer $k$, we say that $g$ is multipodal. We often use the words bipodal and tripodal to mean 2-podal and 3-podal. We say that an $(n+m)$-podal graphon has $(n,m)$ symmetry if $g$ is invariant under permutation of the first $n$ podes and is also invariant under permutation of the remaining $m$ podes. The space of such $(n,m)$-symmetric graphons is only 6-dimensional (or less if $n$ or $m$ is less than 2), with parameters specifying the relative size of the two kinds of podes and the value of $g(x,y)$ on the five (or fewer) equivalence classes of rectangles $I_{i}\times I_{j}$. A graphon with $(2,0)$ symmetry is said to be symmetric bipodal.
In a model with constraints on the densities $\varepsilon,\eta$ of edges and another graph $H$, the phase diagram is the set of all achievable values $(e,t)$ of $(\varepsilon,\eta)$. A phase is a connected open set in the phase diagram where the optimal graphon is unique and is a real analytic function of $(e,t)$. More precisely, for each finite graph $K$ (such as a square or tetrahedron or pentagon), there is a function $\kappa(g)$, similar to $\varepsilon(g)$ and $\tau(g)$, that gives the density of $K$’s in $g$. Within a phase, each such functional applied to the optimal graphon is required to yield an analytic function of $(e,t)$. In practice, analytic dependence of the optimal graphon $g$ on $(e,t)$ is usually proven by describing $g$ (and thus the density of every subgraph $K$) via a finite set of parameters and then showing that these parameters are analytic functions of $(e,t)$. A phase transition occurs where the density of some $K$ is not analytic or is not even defined, such as when the optimal graphon is not unique. Phase transitions have only been shown to occur on boundaries of phases.
1.5. Detailed results
The following theorems give a simple description of what happens near almost all points along the boundary of the Razborov triangle. We prove that the unique optimal graphons in the phases near the boundary are multipodal. The cited theorems in the text include additional estimates on how the podes of the optimal graphons scale as the constraints approach the boundary.
Theorem 1 (Theorem 6).
For each fixed $e<1/2$ and all $t$ sufficiently small, the optimal graphon with edge/triangle densities $(e,t)$ is unique and is symmetric bipodal, with parameters that vary analytically with $(e,t)$.
Theorem 2 (Theorem 7).
Let $n\geq 1$ be an integer. For every $e\in\left(\frac{n}{n+1},\frac{n+1}{n+2}\right)$, with corresponding minimal triangle density $t_{0}$ (depending on $e$), and for all $\Delta t$ sufficiently small, the optimal graphon with edge/triangle densities $(e,t_{0}+\Delta t)$ is unique and $n+2$-podal, with $(n,2)$ symmetry and with parameters that vary analytically with $(e,t)$.
Note that these theorems do not make any claims about what happens exactly over the cusps, i.e. when $e=\frac{n}{n+1}$. When $e=1/2$ ($n=1$), the optimal graphon has long been known to have a symmetric bipodal structure. When $n$ is larger, the optimal graphon is believed to have $(n+1,0)$ symmetry. However, a small neighborhood of each cusp is believed to intersect four(!) different phases, making a precise characterization difficult.
Theorem 3 (Theorem 8).
All of the phases above the scallops proven in Theorems 1 and 2 have unique optimal graphons with distinct symmetries and cannot be analytically continued to one another.
In the notation of Figure 2, Theorem 1 proves that the region just above the flat part of the bottom boundary is part of the $A(2,0)$ phase. Theorem 2 proves the existence of all of the $C(n,2)$ phases, and Theorem 3 shows that these phases are all different.
Theorem 4 (Theorem 9).
For each fixed $e\in(0,1)$ and all $t$ sufficiently close to (but below) $e^{3/2}$, the optimal graphon with edge/triangle densities $(e,t)$ is unique and bipodal, with parameters that vary analytically with $(e,t)$.
That is, the region just below the upper boundary is part of a bipodal phase. There is every reason to believe that this is part of the same bipodal $F(1,1)$ phase that is found just above the ER curve, but this has not yet been proved.
All of these theorems can be viewed as extensions of extremal graph theory. Pikhurko and Razborov’s results [35] determine unique optimal graphons on the boundary of Figure 1 [37]. Theorems 14 describe their unique extensions nearby. This is in contrast to our earlier results that relied on perturbing Erdős-Rényi graphs, or on perturbing known results on the $e=1/2$ line for $t<e^{3}$, all well within the interior of the Razborov triangle.
2. The semi-local “worth” functional
In this section we develop the concept of “worth”. This is a quantity associated with columns of a graphon. We will see that all columns of an optimal graphon must maximize worth. Working with columns, rather than just with the value $g(x,y)$ of the graphon at each point separately, gives us the analytic control needed to prove our main theorems.
We recall the basic theory of Lagrange multipliers. Let $g$ be a graphon that maximizes the Shannon entropy $S(g)$ for particular values $(e,t)$ of $(\varepsilon(g),\tau(g))$. We call an infinitesimal change to the graphon reversible if it is possible to make the opposite change, for which the infinitesimal changes $dS$, $d\varepsilon$, and $d\tau$ are minus those associated with the original change. For instance, increasing the value of a graphon in a portion of $[0,1]^{2}$ from 0.5 to slightly above 0.5 is reversible, since we could just as well lower the value slightly below 0.5. Likewise, expanding the area of the region where $g(x,y)=0.5$ (at the expense of a region where $g(x,y)$ takes on a different value) is reversible, since we could just as well shrink the region. However, increasing the value of a graphon from 0 to a positive number, or increasing the area of a region from 0 to a positive number, are irreversible changes.
Note that the terms reversible and irreversible do not refer to changes in time, as we are not working with any notion of dynamics. Rather, we can visualize infinitesimal changes as tangent vectors to the space of possible graphons. In this heuristic picture, irreversible changes correspond to inward-pointing vectors at the boundary of the space, while reversible changes either refer to points in the interior or are tangent to the boundary.
When considering reversible changes to an optimal graphon, the infinitesimal changes $dS$, $d\varepsilon$ and $d\tau$ to the entropy, edge density and triangle density cannot be linearly independent at $g$. If they were, we could find a reversible change that increased $S$ while leaving $\varepsilon$ and $\tau$ fixed, contradicting the optimality of $g$. There must therefore exist constants $\alpha$ and $\beta$ (depending on $g$) such that, for any reversible infinitesimal change to $g$,
(5) $dS=\alpha d\varepsilon+\frac{\beta}{3}d\tau.$
Equation (5) applies not only at points where $d\varepsilon$ and $d\tau$ are linearly independent, but also on the Erdős-Rényi curve, where $d\varepsilon$ and $d\tau$ are collinear. However, at such points the constants $(\alpha,\beta)$ are not uniquely defined. Rather, each Erdős-Rényi graphon is associated with infinitely many possible values of $(\alpha,\beta)$.
For irreversible changes, the equality is replaced by an inequality: $dS\leq\alpha d\varepsilon+\frac{\beta}{3}d\tau.$ There is no problem with such moves costing us entropy relative to the prevailing exchange rate for $\varepsilon$ and $\tau$, since the opposite move is not possible, but they cannot gain entropy.
This principle is sometimes described in terms of a functional
(6) $F(g)=S(g)-\alpha\varepsilon(g)-\frac{\beta}{3}\tau(g).$
To first order, reversible changes to $g$ cannot change $F$, while irreversible changes can decrease $F$ but cannot increase $F$.
The functionals $S$ and $\varepsilon$ are local, in that there is a contribution from each point $(x,y)\in[0,1]^{2}$ and we integrate the local contributions to get the global quantity. If the triangle density were also local, then $F$ would be the integral of a local density and our variational equations would come from setting the derivative of this density with respect to $g(x,y)$ equal to zero. That is the typical situation when doing calculus of variations, especially in classical and quantum field theory, with the global action being the integral of a local Lagrangian density [5].
However, the triangle density
[…]
is not local. Rather, it involves interactions between the values of $g$ at three different points. To accommodate this complication, we consider (reversible and irreversible) changes to entire columns of $g$ and not just to isolated points. We work up to that in stages, first considering infinitesimal changes to $g$ near a single point, then considering macroscopic changes to $g$ near that point, and finally considering macroscopic changes on an entire column.
If we vary slightly the value of $g(x,y)$ in a neighborhood of a point, then standard methods apply. The functional derivatives of $S$, $\varepsilon$ and $\tau$ with respect to $g(x,y)$ are $H^{\prime}(g(x,y))$, $1$ and $3G(x,y)$, respectively, where
[…]
We conclude that
(9) $H^{\prime}(g(x,y))=\alpha+\beta G(x,y)$
at every point where $g(x,y)$ is continuous. As for points of discontinuity, equation (9) must be understood in a weak sense. If we multiply $H^{\prime}(g(x,y))$ and $\alpha+\beta G(x,y)$ by an arbitrary test function and integrate, the two integrals must be the same.
We can also consider changing the value of $g(x,y)$ by a larger amount, albeit on an infinitesimally small region $R$. Note that this is still an infinitesimal change, in that the resulting (infinitesimal) changes in $\varepsilon$ and $\tau$ can be turned into changes in $S$ by making infinitesimal changes to $g$ elsewhere. The change in $F$ is proportional to the area of $R$ times the change in
(10) $V(x,y)=H(g(x,y))-\alpha g(x,y)-\beta G(x,y)g(x,y).$
The derivative of $V(x,y)$ with respect to $g(x,y)$ is precisely
(11) $H^{\prime}(g(x,y))-\alpha-\beta G(x,y),$
so maxima of $V$ are automatically solutions to (9). Furthermore, the second derivative of $V$ with respect to $g(x,y)$ is always negative, so solutions to (9) always maximize $V$.
An optimal graphon must maximize $V$ at each point, but that local optimization is not sufficient. To account for the non-locality of the $\tau$ function, we define the worth of a function $C:[0,1]\to[0,1]$ to be
[…]
Theorem 5.
If a graphon $g$ maximizes the Shannon entropy $S(g)$ for given values of edge and triangle density, then every column of $g$ maximizes $W$, using the values of $\alpha$ and $\beta$ determined by equation (5) (for instance via equation (9)). In particular, all columns of $g$ must have the same worth.
Proof.
[…]
Using Theorem 5, we can now prove Theorems 1, 2 and 4.
3. Proof of Theorem 1
We will prove a slightly more quantitative version of Theorem 1:
Theorem 6.
For each fixed $e<1/2$ and all $t$ sufficiently small, the optimal graphon with edge/triangle densities $(e,t)$ is unique and is symmetric bipodal, with parameters that vary analytically with $(e,t)$. As $t\to 0$, the increase $\Delta\mathbb{B}=\mathbb{B}(e,t)-\mathbb{B}(e,0)$ in the Boltzmann entropy scales as $t\ln(1/t)$ and the Lagrange multiplier $\beta$ scales as $\ln(1/t)$.
3.1. Strategy
The proofs of Theorems 6, 7 and 9 all follow the same general outline. We will present the proof of Theorem 6 in full detail. The subsequent proofs of Theorems 7 and 9 will be somewhat abbreviated, concentrating on what is different in those cases.
Using the fact that the unique entropy maximizing graphon $g_{0}$ at $(e,0)$ is symmetric bipodal, we show that the optimizing graphons at points near the boundary have the same general structure away from a small exceptional set. Specifically, we partition the unit interval into subsets $I_{1}$, $I_{2}$ and $I_{3}$ such that the columns $g_{x}$ of the optimal graphon are $L^{2}$-close to the columns of the first pode of $g_{0}$ when $x\in I_{1}$ and are $L^{2}$-close to the columns of the second pode of $g_{0}$ when $x\in I_{2}$, and where $I_{3}$ has small measure. At this stage, we do not have any control over $g_{x}$ when $x\in I_{3}$.
Knowing the columns $g_{x}$ when $x\in I_{1}\cup I_{2}$ (to within a small error in $L^{2}$) gives us pointwise control of the function $G(x,y)$ on $(I_{1}\cup I_{2})\times(I_{1}\cup I_{2})$. The Euler-Lagrange equations (9) then give us pointwise estimates of $g(x,y)$ in each of the four main rectangles.
[…]
Using the pointwise equations (9), we bound the maximum fluctuation of $g(x,y)$ in each rectangle $I_{i}\times I_{j}$ by a multiple of the maximum fluctuation in a neighboring rectangle. Combining these results, the maximum fluctuation in each rectangle is bounded by a small multiple of itself, and so must be zero. That is, our optimal graphon must be exactly bipodal.
The space of bipodal graphons with given values of $(e,t)$ is only 2-dimensional. Using ordinary calculus, we determine that the entropy is maximized when the graphon is symmetric.
3.2. Defining approximate podes
There is a unique entropy maximizer $g_{0}$ at $(e,0)$ on the bottom boundary of the Razborov triangle (up to measure-preserving transformations of the unit interval, as usual). This graphon is symmetric bipodal, taking values 0 on the diagonal blocks and $2e$ on the off-diagonal blocks. As we approach the bottom boundary of the Razborov triangle, we claim that any sequence $\{g_{i}\}$ of entropy maximizers must converge (after appropriate measure-preserving transformations) to $g_{0}$ in $L^{2}$.
To see this, we invoke the compactness of the space of reduced graphons in the cut metric. A subsequence must converge to a limit $g_{\infty}$ in the cut metric. By the semicontinuity of the entropy and the fact that the limit of the entropies of the entropy maximizers is at least $S(g_{0})$, we must have $S(g_{\infty})\geq S(g_{0})$. But $g_{0}$ is the unique entropy maximizer at $(e,0)$, so $g_{\infty}=g_{0}$.
The entire sequence $\{g_{i}\}$, and not just a subsequence, must converge to $g_{0}$. If it did not, we could find a subsequence where all points were bounded away from $g_{0}$ in the cut metric. Applying the previous argument to this subsequence would then yield a contradiction.
The entropy $S(g)$ is continuous in $L^{2}$ but only semicontinuous in the cut metric. Limits in the cut metric that are not limits in $L^{2}$ are associated with jumps in the entropy. Since $S(g_{0})=\lim S(g_{i})$, the convergence of $g_{i}$ to $g_{0}$ must occur in $L^{2}$ as well as in the cut metric. That is, for every $\epsilon>0$ there is a $\delta>0$ such that, for all $t<\delta$ and all optimal graphons $g$ at $(e,t)$, $\|g-g_{0}\|_{L^{2}}<\epsilon$. (Note that we have not assumed that the optimal graphon $g$ is unique. That will be proven in due course.)
Let $g$ be such an optimal graphon for a particular value of $(e,t)$. Then
[…]
so
[…]
except on a set of $x$’s of measure $\epsilon$ or less. Call that exceptional set $I_{3}$. Let $I_{1}$ and $I_{2}$ be the intersection of $I_{3}^{c}$ with $[0,1/2]$ and $[1/2,1]$, respectively. That is, we have broken the unit interval into three pieces $I_{1}$, $I_{2}$, $I_{3}$, such that:
  • For all $x\in I_{1}$, $\|g_{x}-C_{1}\|<\sqrt{\epsilon}$.
  • For all $x\in I_{2}$, $\|g_{x}-C_{2}\|<\sqrt{\epsilon}$.
  • When $x\in I_{3}$ we do not yet have any estimates on $g_{x}$.
Here $C_{1}$ and $C_{2}$ are the columns of $g_{0}$ on the two podes, namely $2e$ times the indicator functions of $[1/2,1]$ and $[0,1/2]$, respectively. We will refer to the sets $I_{1}$, $I_{2}$ and $I_{3}$ as podes, even though we are not assuming that the graphon $g$ is exactly tripodal.
3.3. Variational equations
We examine the terms in our variational equations (9). The first two derivatives of the function $H$ are
(15) $H^{\prime}(u)=\ln(1-u)-\ln(u),\qquad H^{\prime\prime}(u)=-\left(\frac{1}{u}+\frac{1}{1-u}\right).$
The quantity $G(x,y)$ is the $L^{2}$-inner product of $g_{x}$ and $g_{y}$, which we denote $\langle g_{x}|g_{y}\rangle$. That is,
[…]
If $x$ and $y$ are both in $I_{1}$, or both in $I_{2}$, then $G(x,y)=2e^{2}+O(\sqrt{\epsilon})$. If one is in $I_{1}$ and the other is in $I_{2}$, then $G(x,y)=O(\sqrt{\epsilon})$. If either or both are in $I_{3}$, then our estimates do not apply.
For $(x,y)\in I_{1}\times I_{1}$ or $I_{2}\times I_{2}$, we have
(17) $H^{\prime}(g(x,y))=\alpha+2\beta e^{2}(1+O(\sqrt{\epsilon}))=2\beta e^{2}(1+O(\sqrt{\epsilon})),$
so
(18) $g(x,y)=\exp(-2e^{2}\beta(1+O(\sqrt{\epsilon})).$
(Since $\beta$ is divergent as $t\to 0$ but $\alpha$ is not, we can absorb $\alpha$ into the $O(\beta\sqrt{\epsilon})$ error.) This means that the contribution of $g(x,y)$ in $I_{1}\times I_{1}$ or $I_{2}\times I_{2}$ to $\beta G$ goes as $\beta$ times a negative exponential in $\beta$, and thus has an extremely small effect on the value of $g(x,y)$ in $I_{1}\times I_{2}$ or $I_{2}\times I_{1}$.
However, we cannot yet precisely estimate $g(x,y)$ in those regions because $G(x,y)=\langle g_{x}|g_{y}\rangle$ also gets a contribution, potentially of order $\epsilon$, from $z\in I_{3}$.
3.4. Maximizing worth and eliminating $I_{3}$
Let $C:[0,1]\to[0,1]$ be a function whose worth we wish to estimate. Let
[…]
That is, $a$ and $b$ are the average values of $C$ on $[0,1/2]$ and $[1/2,1]$.
We now consider the three expressions that contribute to $W(C)$:
[…]
Up to small deviations due to the differences between $g$ and $g_{0}$, maximizing $W(C)$ amounts to taking $C(y)$ to be constant on $[0,1/2]$ and constant on $[1/2,1]$ and choosing $a$ and $b$ to maximize
(20) $H(a)+H(b)-\alpha(a+b)-e\beta ab.$
Setting the derivatives of (20) to zero gives the equations
(21) $H^{\prime}(a)=\alpha+\beta eb,\qquad H^{\prime}(b)=\alpha+\beta ea.$
Since there is a worth-maximizer with $a$ close to 0 and $b$ close to $2e$ (namely any column with $x\in I_{1}$), and another worth-maximizer with $a$ close to $2e$ and $b$ close to 0, $\alpha$ must be close to $H^{\prime}(2e)$, while $\beta$ is large and positive.
If $a$ is substantially nonzero (say, bigger than $1/\sqrt{\beta}$), then $e\beta a$ is gigantic and $b$ is extremely close to zero, being $O(\exp(-\sqrt{\beta}))$. This makes $e\beta b$ tiny so $H^{\prime}(a)\approx\alpha\approx H^{\prime}(2e)$ and $a\approx 2e$. Similarly, if $b$ is substantially nonzero then $a$ is tiny and $b\approx 2e$. In both those cases, $W(C)\approx\frac{1}{2}H(2e)-eH^{\prime}(2e)=-\frac{1}{2}\ln(1-2e)$. The third possibility is that $a$ and $b$ are both tiny, but in that case $W(C)\approx 0$, which is strictly less than $-\frac{1}{2}\ln(1-2e)$.
The upshot is that there are three stationary points of (20) but only two maxima, one that resembles $g_{x}$ for $x\in I_{1}$ and one that resembles $g_{x}$ for $x\in I_{2}$. Since every column $g_{x}$ with $x\in I_{3}$ must be a worth-maximizer, and since every worth-maximizer must come close to maximizing (20), every column $g_{x}$ with $x\in I_{3}$ resembles $g_{x}$ for $x\in I_{1}$ or $I_{2}$. We can then reassign the points of $I_{3}$ to $I_{1}$ or $I_{2}$ depending on the nature of $g_{x}$.
3.5. Exact bipodality
Our next step is to upgrade our $L^{2}$ estimates on the forms of the different columns into pointwise estimates. Thanks to each column of $g$ being $L^{2}$-close to a column of $g_{0}$, the function $G(x,y)=\langle g_{x}|g_{y}\rangle$ is close to $2e^{2}$ on $I_{1}\times I_{1}$ and on $I_{2}\times I_{2}$. By (9), this forces $g(x,y)$ to be exponentially small (in $\beta$) in these quadrants. This in turn makes $G(x,y)$ exponentially small on $I_{1}\times I_{2}$ and $I_{2}\times I_{1}$, which means that $H^{\prime}(g)$ is exponentially close to $\alpha$ in these quadrants, and therefore that $g(x,y)$ is pointwise close to constant in these rectangles.
We next show that the optimal graphon $g$ is exactly constant on each of those rectangles. Let $A$, $B$, and $D$ be the average values of $g(x,y)$ on $I_{1}\times I_{1}$, $I_{2}\times I_{2}$ and $I_{1}\times I_{2}$, respectively. Let $\Delta_{A}$, $\Delta_{B}$ and $\Delta_{D}$ be the difference between the maximum and minimum values of $g(x,y)$ on those rectangles. Let $c$ be the width of $I_{1}$.
On $I_{1}\times I_{1}$, the quantity $G(x,y)$ is bounded below by
(22) $c(A-\Delta_{A})^{2}+(1-c)(D-\Delta_{D})^{2}$
and bounded above by
(23) $c(A+\Delta_{A})^{2}+(1-c)(D+\Delta_{D})^{2}.$
The difference between these two expressions is $4cA\Delta_{A}+4(1-c)D\Delta_{D}$.
All points satisfy the variational equations
(24) $H^{\prime}(g(x,y))=\alpha+\beta G(x,y).$
Subtracting this equation at the smallest value of $G(x,y)$ from that at the largest value, applying the mean value theorem to the left hand side, and applying our bounds on the variation in $G(x,y)$, we obtain
(25) $-H^{\prime\prime}(A_{0})\Delta_{A}\leq 4Ac\beta\Delta_{A}+4D(1-c)\beta\Delta_{D},$
where $A_{0}$ is some number between $A+\Delta_{A}$ and $A-\Delta_{A}$. A little algebra then shows that
(26) $\Delta_{A}\leq\frac{4D(1-c)\beta}{-H^{\prime\prime}(A_{0})-4Ac\beta}\Delta_{D}\leq\frac{3D\beta}{-H^{\prime\prime}(A)}\Delta_{D},$
where we have used the difference between 3 and $4(1-c)\approx 2$ to cover for simplifying the denominator and replacing $A_{0}$ with $A$. A similar analysis on $I_{2}\times I_{2}$ shows that
(27) $\Delta_{B}\leq\frac{3D\beta}{-H^{\prime\prime}(B)}\Delta_{D}.$
Meanwhile, on $I_{1}\times I_{2}$, $G(x,y)$ is bounded above and below by
(28) $c(A\pm\Delta_{A})(D\pm\Delta_{D})+(1-c)(B\pm\Delta_{B})(D\pm\Delta_{D}),$
where the plus signs give an upper bound and the minus signs give a lower bound. The difference between the upper and lower bounds is
(29) $2\beta[(cA+(1-c)B)\Delta_{D}+(c\Delta_{A}+(1-c)\Delta_{B})D].$
This implies that
(30) $-H^{\prime\prime}(D_{0})\Delta D\leq 2\beta(cA+(1-c)B)\Delta_{D}+2\beta D(c\Delta_{A}+(1-c)\Delta_{B}).$
A little algebra then gives
(31) $\Delta_{D}$$\leq$ $\frac{2\beta D(C\Delta A+(1-c)\Delta B)}{-H^{\prime\prime}(D_{0})-2\beta(cA+(1-c)B}$
(32) $\leq$ $\frac{3\beta D(\Delta_{A}+\Delta_{B})}{-2H^{\prime\prime}(D)}$
(33) $\leq$ $\frac{9\beta^{2}D^{2}}{-2H^{\prime\prime}(D)}\left(\frac{-1}{H^{\prime\prime}(A)}+\frac{-1}{H^{\prime\prime}(B)}\right)\Delta_{D}.$
Now recall that $A$ and $B$ are exponentially small in $\beta$ and that
(34) $\frac{-1}{H^{\prime\prime}(A)}=A(1-A)<A\qquad\hbox{and}\qquad\frac{-1}{H^{\prime\prime}(B)}=B(1-B)<B.$
The coefficient of $\Delta_{D}$ on the right hand side of the last line goes to zero roughly as $\beta^{2}\exp(-2e^{2}\beta)$ as $t\to 0$ and $\beta\to\infty$. Once the coefficient is less than one, the only solution is $\Delta_{D}=0$, which then implies that $\Delta_{A}=0$ and $\Delta_{B}=0$. In other words, our optimal graphon is exactly bipodal.
3.6. Symmetric bipodality
All that remains is showing that the best bipodal graphon is symmetric, with pode sizes $\frac{1}{2}$ and $\frac{1}{2}$ and with $A=B$. This requires extensive calculations but no sophisticated analysis. Ultimately, it is just a (grungy) problem in multivariable calculus.
For each triple $(e,t,c)$ we consider the bipodal graphon that maximizes the entropy, subject to the constraints that the edge and triangle densities are $(e,t)$ and that the first pode has width $c$. Let $S(e,t,c)$ be the entropy of this optimal graphon. We must show that this entropy is maximized at $c=1/2$. Note that this function is analytic in $c$ for fixed $(e,t)$, insofar as the parameters are determined by analytic Euler-Lagrange equations, and is even in $\Delta c:=c-\frac{1}{2}$.
When $t=0$, the function is easy to compute. The graphon must be zero on $I_{1}\times I_{1}$ and $I_{2}\times I_{2}$ and take on the constant value $\frac{e}{2c(1-c)}=\frac{2e}{1-4\Delta c^{2}}$ on $I_{1}\times I_{2}$. The entropy is then
(35) $S(e,0,c)$ $=$ $\frac{1}{2}(1-4\Delta c^{2})H\left(\frac{2e}{1-4\Delta c^{2}}\right)$
(36) $=$ $S(e,0,1/2)+2\ln(1-2e)\Delta c^{2}+O(\Delta c^{4}).$
That is, there is an entropy cost proportional to $\Delta c^{2}$ associated with having $\Delta c\neq 0$.
Now consider the effect of having $t$ nonzero. Having the graphon nonzero on $I_{1}\times I_{1}$ and $I_{2}\times I_{2}$ provides additional entropy of order $t\ln(1/t)$. Shifting the value of the graphon on $I_{1}\times I_{2}$ by an $O(t)$ amount changes the entropy by an additional $O(t)$, but since this is small compared to $t\ln(1/t)$, $S(e,t,c)-S(e,0,c)$ is still $O(t\ln(1/t))$. In order to overcome the $-2\ln(1-2e)\Delta c^{2}$ cost, we must have $\Delta c=O(\sqrt{t\ln(1/t)})$. Since $t\sim\exp(-2e^{2}\beta)$, this means that $\Delta c$ must be exponentially small in $\beta$ and in particular that $\beta\Delta c$ is a small parameter.
We now compute the quantity $G(x,y)$ in each rectangle and look at the Euler-Lagrange equations for a particular value of $\beta$:
(37) $H^{\prime}(A)$ $=$ $\alpha+\frac{\beta}{2}(A^{2}+D^{2})-\beta\Delta c(D^{2}-A^{2})$
(38) $\approx$ $\alpha+\frac{\beta}{2}D^{2}-\beta\Delta cD^{2},$
(39) $H^{\prime}(B)$ $=$ $\alpha+\frac{\beta}{2}(B^{2}+D^{2})+\beta\Delta c(D^{2}-B^{2})$
(40) $\approx$ $\alpha+\frac{\beta}{2}D^{2}+\beta\Delta cD^{2},$
(41) $H^{\prime}(D)$ $=$ $\alpha+\frac{\beta D}{2}(A+B+2\Delta c(A-B))$
(42) $\approx$ $\alpha,$
where in our approximations we use the fact that $A$ and $B$ are exponentially small in $\beta$. Since $D\approx 2e$, this makes $\alpha\approx H^{\prime}(2e)$. The terms proportional to $\Delta c$ serve to multiply $A$ by a factor of $\exp(-4e^{2}\beta\Delta c)\approx 1-4e^{2}\beta\Delta c$ and to multiply $B$ by a factor of $\exp(4e^{2}\beta\Delta c)\approx 1+4e^{2}\beta\Delta c$. These changes in the values of $A$ and $B$ (relative to their values when $\Delta c=0$) slightly change the triangle density for a given value of $\beta$, but only by a fraction $O(\beta\Delta c^{2})$. Likewise, the contribution to the entropy of the $I_{1}\times I_{1}$ and $I_{2}\times I_{2}$ squares changes by a fraction $O(\beta\Delta c^{2})$. However, that entropy is only $O(t\ln(1/t))$, so we are dealing with an expression that is
(43) $O(\beta t\ln(1/t)\Delta c^{2})=O(t(\ln(1/t)^{2})\Delta c^{2}),$
since $\beta=O(\ln(1/t))$. This possible entropy gain from having $\Delta c\neq 0$ is much smaller than the $-2\ln(1-2e)\Delta c^{2}$ cost, so the optimal value of $\Delta c$ is exactly zero. That is, we must have $c=1/2$.
When $c=1/2$, two of the Euler-Lagrange equations read:
(44) $H^{\prime}(A)$ $=$ $\alpha+\beta(A^{2}+D^{2})/2,$
(45) $H^{\prime}(B)$ $=$ $\alpha+\beta(B^{2}+D^{2})/2.$
If $A>B$, then the right hand side of the first equation is greater than that of the second, so $H^{\prime}(A)>H^{\prime}(B)$. But that is a contradiction, since $H^{\prime}(u)=\ln(1-u)-\ln(u)$ is a decreasing function of $u$. Likewise, we cannot have $A<B$. So $A$ and $B$ must be equal, making our optimal graphon symmetric bipodal.
The parameters of a symmetric bipodal graphon are uniquely (and analytically) determined by $(e,t)$.
We also consider how various quantities scale as $t\to 0$. After setting $c=1/2$ and $B=A$, a direct calculation shows that
(46) $t=\frac{3}{4}AD^{2}+\frac{1}{4}A^{3},$
so
(47) $A=\frac{4t}{3D^{2}}+O(t^{3})=\frac{t}{3e^{2}}+O(t^{2}),$
where we have used the fact that $D=2e-A$. Since $A$ was exponentially small in $\beta$, $\beta$ must scale as $\ln(1/t)$. The entropy is
(48) $\frac{1}{2}(H(A)+H(2e-A))=\frac{1}{2}H(2e)-\frac{1}{2}A\ln(A)+O(A),$
so $S(g)-\frac{1}{2}H(2e)$ scales as $t\ln(1/t)$.
The Boltzmann entropy $\mathbb{B}(e,t)$ is equal to the Shannon entropy $S(g)$ of the optimal graphon at $(e,t)$, so $\Delta\mathbb{B}:=\mathbb{B}(e,t)-\mathbb{B}(e,0)=S(g)-\frac{1}{2}H(2e)$. Since $A\approx(t/3e^{2})$ and $A\approx\exp(-2e^{2}\beta)$, $\beta\approx\frac{-1}{2e^{2}}\ln(t/3e^{2})\sim\ln(1/t)$. ∎
4. Proof of Theorem 2
As in the last section, we will prove a slightly extended version of Theorem 2:
Theorem 7.
Let $n\geq 1$ be an integer. For every $e\in\left(\frac{n}{n+1},\frac{n+1}{n+2}\right)$, with corresponding minimal triangle density $t_{0}$ (depending on $e$), and for all $\Delta t$ sufficiently small, the optimal graphon with edge/triangle densities $(e,t_{0}+\Delta t)$ is unique and $n+2$-podal, with $(n,2)$ symmetry. Asymptotically, $\Delta\mathbb{B}=\mathbb{B}(e,t)-\mathbb{B}(e,t_{0})$ scales as $\sqrt{\Delta t}$. and the Lagrange multiplier $\beta$ scales as $1/\sqrt{\Delta t}$. In the optimal graphon, the diagonal entries are all $\exp(-\Theta(\beta))$ and, except for the $(n+1,n+2)$ and $(n+2,n+1)$ entries, the off-diagonal entries are all $1-\exp(-\Theta(\beta))$.
Proof.
The proof of Theorem 7 (and therefore Theorem 2) follows the same script as the proof of Theorem 6, namely
  1. (1)
    Using the proximity of an entropy-maximizing graphon $g$ at $(e,t)$ to the unique entropy-maximizing graphon $g_{0}$ at $(e,t_{0})$ to define approximate podes $I_{1},\ldots,I_{n+3}$ where the columns with $x\in I_{j}$ with $j\leq n+2$ are $L^{2}$-close to the corresponding columns of $g_{0}$, and where the exceptional set $I_{n+3}$ is small.
  2. (2)
    Using the Euler-Lagrange equations to show that all of the graphon values are exponentially close to 0 or 1, except on $I_{n+1}\times I_{n+2}$, $I_{n+2}\times I_{n+1}$, or when one of the coordinates is in $I_{n+3}$.
  3. (3)
    Showing that the only possible worth-maximizing columns are small perturbations of the columns of $g_{0}$, thus allowing us to reassign the points of $I_{n+3}$ to the other podes.
  4. (4)
    Bounding the variation in $g(x,y)$ in each rectangle $I_{i}\times I_{j}$ by a small multiple of the variation in other rectangles. Combining estimates, this shows that the variation in each rectangle is bounded by a small multiple of itself, and must therefore be zero.
  5. (5)
    Analyzing the finite-dimensional space of $(n+2)$-podal graphons near $g_{0}$ and determining that the best one has $(n,2)$ symmetry. We then determine how $S$, $\beta$, and various entries of the optimal graphon scale with $\Delta t$.
There is one important difference between the situation of Theorem 6 and that of Theorem 7. The additional podes that appear on the scallops provide an additional, and more efficient, means of generating entropy at the expense of added triangles. As a result, $\Delta\mathbb{B}$ scales as $\sqrt{\Delta t}$ rather than $\Delta t\ln(1/\Delta t)$. Before getting into the details of the proof, we explain how this works, starting near the first scallop, with $e\in(\frac{1}{2},\frac{2}{3})$.
Consider tripodal graphons of the form shown in Figure 4. The total edge density is
(49) $e=2c(1-c)+\frac{1}{2}p(1-c)^{2},$
so we must have
(50) $p=\frac{2(e-2c(1-c))}{(1-c)^{2}}.$
The triangle density is
(51) $t=\frac{3}{2}pc(1-c)^{2}=3c(e-2c(1-c))=3ec-6c^{2}+6c^{3}.$
Taking derivatives, we see that
(52) $\frac{dt}{dc}=3(6c^{2}-4c+e)\hbox{ and }\frac{d^{2}t}{dc^{2}}=36c-12.$
The first derivative is zero when
(53) $c=\frac{1}{3}\left(1+\sqrt{1-\frac{3e}{2}}\right).$
Since $d^{2}t/dc^{2}$ is always positive, this gives the minimum triangle density among graphons of this kind. In fact, it minimizes $t$ among all possible graphons [35] and is the unique optimal graphon with densities $(e,t_{0})$ [37].
Figure 4. A tripodal graphon of the form seen on the first scallop
Now imagine varying $c$ and $p$ while preserving the structure of Figure 4. The entropy of the graphon displayed in Figure 4 is
(54) $S=\frac{1}{2}(1-c)^{2}H(p),$
where $p$ is given by equation (50). A little algebra then gives
(55) $p=4-4(1-c)^{-1}+2e(1-c)^{-2},$
so
(56) $\frac{dp}{dc}=4e(1-c)^{-3}-4(1-c)^{-2}=\frac{4(e-(1-c))}{(1-c)^{3}}.$
We then compute
(57) $\frac{dS}{dc}$ $=$ $-(1-c)H(p)+\frac{1}{2}(1-c)^{2}H^{\prime}(p)\frac{dp}{dc}$
(58) $=$ $-(1-c)H(p)+\frac{2H^{\prime}(p)(e+c-1)}{1-c}$
(59) $=$ $-(1-c)H(p)+\frac{H^{\prime}(p)}{1-c}(p(1-c)^{2}+6c-4c^{2}-2)$
(60) $=$ $(1-c)(pH^{\prime}(p)-H(p))+(4c-2)H^{\prime}(p)$
(61) $=$ $(1-c)\ln(1-p)+4c-2)(\ln(1-p)-\ln(p))$
(62) $=$ $(3c-1)\ln(1-p)+2(1-2c)\ln(p).$
Since $c$ is between $\frac{1}{2}$ and $\frac{2}{3}$, the coefficients of $\ln(1-p)$ and $\ln(p)$ are both positive, making $\frac{dS}{dc}$ negative. We can increase the entropy to first order by decreasing $c$. That only increases the triangle count to second order in $\Delta c$, so we have achieved an entropy increase that scales as the square root of $\Delta t$.
The situation is similar near the other scallops. There is a family of graphons parametrized by the size $c$ of each of the $n$ identical podes, as in Figure 5 for $n=3$. There is a value $c_{0}$ that minimizes the triangle density, but $dS/dc$ is not zero at $c=c_{0}$. Instead, the calculation shown in the next paragraph shows that $dS/dc$ is negative for all relevant values of $c$. As a result, we can increase $S$ to first order in $\Delta c$ by decreasing $c$ while only increasing $t$ to second order, so $\Delta S\sim\sqrt{\Delta t}$.
Figure 5. A multipodal graphon of the form seen on the scallops, in this case with $n=3$
For a general value of $n$, the graphon on the $n$-th scallop is 1 everywhere except on the diagonal blocks and on the two off-diagonal blocks in the upper right corner. The edge density is
(63) $e=1-nc^{2}-(1-nc)^{2}+\frac{(1-nc)^{2}}{2}p.$
This means that
(64) $p$ $=$ $\frac{2}{(1-nc)^{2}}\left(e+nc^{2}-1+(1-nc)^{2}\right)$
(65) $=$ $2\left(e-\frac{n-1}{n}\right)(1-nc)^{-2}-\frac{4}{n}(1-nc)^{-1}+\frac{2(n+1)}{n}.$
Taking a derivative with respect to $c$ is then easy:
(66) $\frac{dp}{dc}=\frac{4n}{(1-nc)^{3}}(e+c-1).$
The entropy is $S=\frac{(1-nc)^{2}}{2}H(p)$ and derivative of $S$ with respect to $c$ is
(67) $\frac{dS}{dc}$ $=$ $-n(1-nc)H(p)+\frac{1}{2}(1-nc)^{2}H^{\prime}(p)\frac{dp}{dc}$
(68) $=$ $-n(1-nc)H(p)+\frac{2nH^{\prime}(p)(e+c-1)}{1-nc}$
(69) $=$ $-n(1-nc)H(p)+nH^{\prime}(p)((n+1)c-1+p)$
(70) $=$ $n(1-nc)(pH^{\prime}(p)-H(p))+2n((n+1)c-1)H^{\prime}(p)$
(71) $=$ $-n(1-nc)\ln(1-p)+2n((n+1)c-1)(\ln(1-p)-\ln(p))$
(72) $=$ $n((n+2)c-1)\ln(1-p)+2n(1-(n+1)c)\ln(p).$
Since $c$ is between $\frac{1}{n+2}$ and $\frac{1}{n+1}$, the coefficients of $\ln(p)$ and $\ln(1-p)$ are both positive, making each term negative, so $\frac{ds}{dc}<0$, as claimed.
4.1. Defining approximate podes
We now turn to the details of the proof. As usual, let $g_{0}$ be the unique entropy maximizer $g_{0}$ at $(e,t_{0})$. This graphon takes the form shown in Figure 5, with
(73) $c=c_{0}=\frac{1+\sqrt{1-\frac{n+2}{n+1}e}}{n+2},$
which is the value of $c$ that minimizes
(74) $t=n(n+1)(n+2)c^{3}-3n(n+1)c^{2}+3nec.$
As we approach the scallop, any sequence of entropy maximizers must converge to $g_{0}$ in $L^{2}$ by exactly the same argument as in the proof of Theorem 6.
As before, we pick a sufficiently small value of $\epsilon$ and consider values of $t$ small enough that $\|g-g_{0}\|_{L^{2}}<\epsilon$ for each optimal graphon $g$. Let $I_{1}$, $I_{2}$, $\ldots$, $I_{n+2}$ be the subsets of the podes of $g_{0}$ for which $g_{x}$ lies within $\sqrt{\epsilon}$ in $L^{2}$ of the corresponding column of $g_{0}$ and let $I_{n+3}$ be the exceptional set where $g_{x}$ is not close to the corresponding column of $g_{0}$. Note that $I_{n+3}$ may contain points $x$ where $g_{x}$ is close to a different column of $g_{0}$. Those points will soon be reassigned.
4.2. Variational equations
Next we need to compute $G(x,y)$ in different cases. Let $G_{i,j}$ denote a typical value of $G(x,y)$ when $x\in I_{i}$ and $y\in I_{j}$. We can estimate these quantities to within $O(\sqrt{\epsilon})$ using the columns of $g_{0}$. Thanks to our $(n,2)$ symmetry, there are only five different numbers to compute, namely $G_{1,1}$, $G_{1,2}$, $G_{1,{n+1}}$, $G_{n+1,n+1}$ and $G_{n+1,n+2}$. The results are
(75) $G_{1,1}$$\approx$ $1-c,$
(76) $G_{1,2}$$\approx$ $1-2c,$
(77) $G_{1,n+1}$$\approx$ $(n-1)c+\frac{1-nc}{2}p,$
(78) $G_{n+1,n+1}$$\approx$ $nc+\frac{1-nc}{2}p^{2},$
(79) $G_{n+1,n+2}$$\approx$ $nc,$
where $\approx$” means “equal to within $O(\sqrt{\epsilon})$”. Note that $G_{1,1}$ and $G_{n+1,n+1}$ are greater than $G_{n+1,n+2}$ by amounts that are $\Omega(1)$ as $t\to 0$ while $G_{1,2}$, and $G_{1,n+1}$ are less than $G_{n+1,n+2}$ by amounts that are $\Omega(1)$. Multiplying by $\beta$ and adding $\alpha$, and using the fact that
(80) $H^{\prime}(p)=\alpha+\beta G_{n+1,n+2}=(\alpha+nc)+O(\beta\sqrt{\alpha}).$
we get that $H^{\prime}(g(x,y))$ is $\Omega(\beta)$ on the diagonal rectangles that do not involve $I_{n+3}$ and is $-\Omega(\beta)$ on the off-diagonal blocks that do not involve $I_{n+3}$, with the exception of $I_{n+1}\times I_{n+2}$ and $I_{n+2}\times I_{n+1}$. This implies that $g(x,y)$ is exponentially small on the diagonal blocks, exponentially close to 1 on all but two of the off-diagonal blocks, and of course is close to $p$ on $I_{n+1}\times I_{n_{2}}$ and $I_{n+2}\times I_{n+1}$.
4.3. Maximizing worth
Let $C:[0,1]\to[0,1]$ be a function whose worth we aim to maximize. For each $i=1,2,\ldots,n+2$, let $a_{n+2}$ be the average values of $C(y)$ for $y\in I_{i}$. Of the terms contributing to $W(C)$, the entropy term is bounded above by $c\sum_{i=1}^{n}H(a_{i})+\frac{1-nc}{2}(H(a_{n+1})+H(a_{n+2}))$, since fluctuations in $C$ within each pode can only decrease the entropy. The edge density term is $-\alpha(c\sum_{i=1}^{n}a_{i}+\frac{1-nc}{2}(a_{n+1}+a_{n+2}))$.
The most important term comes from triangles. To within the accuracy of our approximation that $g(y,z)$ is constant on each rectangle, it is the quadratic function
(81) $-\frac{\beta}{2}\sum_{i,j=1}^{n+2}M_{ij}a_{i}a_{j},$
where $M_{ij}$ is the integral of $g(y,z)$ over $I_{i}\times I_{j}$.
If we are at a maximum of $W$, then the gradient of $W$ must be zero and the Hessian must be negative semi-definite. The Hessian of $W$ with respect to the variables $\{a_{i}\}$ is precisely $-\beta M$ plus diagonal terms proportional to $H^{\prime\prime}(a_{i})$. The matrix $M$ is (nearly) zero on the diagonal, with all of the off-diagonal terms being close to 1 or $p$, and so has eigenvalues of both signs. The only way for the Hessian to be negative semi-definite is for all but one of entries $H^{\prime\prime}(a_{i})$ to be at least of order $\beta$. In other words, all columns that maximize $W$ must have every entry but one (or every entry) approximately equal to 0 or 1. In terms of $G$, for $y$ in any pode but one, $|G(x,y)-nc|$ must exceed $\Theta(1/\beta)$.
We now examine the possibilities.
  • If $G(x,y)\approx nc$ for $y\in I_{1}$, then $c\sum_{i=2}^{n}a_{i}+\frac{1-nc}{2}(a_{n+1}+a+{n+2})\approx nc$. But that is impossible if each $a_{i}$ (other than $a_{1}$) is equal to 0 or 1. Contradiction. Likewise, it is not possible to have $G(x,y)\approx nc$ for $y\in I_{2},\ldots,I_{n}$. The first $n$ variables $a_{i}$ are all either pegged to 0 or to 1.
  • If two or more of the variables $a_{1},\ldots,a_{n}$ are pegged to 0, then $G(x,y)<nc$ for all $y$, so $g(x,y)\approx 1$ for all $y$, which is a contradiction. Thus either one or none of the first $n$ $a_{i}$’s is pegged to 0 and the rest are pegged to 1.
  • If exactly one of these variables is pegged to 0, then for $y\in I_{n+1}$ or $y\in I_{n+2}$ we have $G(x,y)\leq p\frac{1-c}{2}+(n-1)c<nc$, so $a_{n+1}$ and $a_{n+2}$ are pegged to 1. In other words, our column is just like the columns when $x$ is in one of the first $n$ podes.
  • If all of the variables $a_{1},\ldots,a_{n}$ are pegged to 1, then we examine $a_{n+1}$ and $a_{n+2}$. Neither one is pegged to 1, since $G(x,y)$ is at least $nc$ for $y$ in either $I_{n+1}$ or $I_{n+2}$. They cannot both be pegged to 0, since that would make $G(x,y)=nc$ in both podes, meaning that the values are not pegged and the Hessian is not negative-definite. Thus one value must be pegged to 0 while the other is intermediate between 0 and 1. The closeness of all but one $a_{i}$ to 0 or 1 gives us the same equation for the remaining $a_{i}$ as satisfied by the actual columns of $I_{n+1}$ or $I_{n+1}$, implying that the final $a_{i}$ must be close to $p$. That is, $C$ is close to the actual columns when $x\in I_{n+1}$ or $I_{n+2}$.
The upshot is that all worth-maximizers are already $L^{2}$-close to columns of $g_{0}$. Since each column is a worth-maximizer, we can reassign all of the points of $I_{n+3}$ to other podes. Note that this reassignment can result in the podes of $g$ having sizes that are slightly different from those of the corresponding podes of $g_{0}$.
Controlling the columns to within a small $L^{2}$ error gives us pointwise control over $G(x,y)=\langle g_{x}|g_{y}\rangle$. On all rectangles except for $I_{n+1}\times I_{n+2}$ and $I_{n+2}\times I_{n+1}$, this forces $g(x,y)$ to be exponentially close to 0 or 1. This makes $G(x,y)$ exponentially close to constant on $I_{n+1}\times I_{n+2}$ and $I_{n+2}\times I_{n+1}$ and so makes $g(x,y)$ exponentially close to a constant (that is close to $p$, but not necessarily exponentially close) on these rectangles.
4.4. Exact multipodality
So far we have shown that an optimal graphon has to be approximately multipodal. There are $n$ podes $I_{1},\ldots,I_{n}$ of width close to
(82) $c=\frac{1+\sqrt{1-\frac{n+2}{n+1}e}}{n+2}$
and two podes of width close to $\frac{1-nc}{2}$. The graphon is exponentially close to 0 on the diagonal blocks, exponentially close to 1 on all of the off-diagonal blocks but two, and close to $p$ on $I_{n+1}\times I_{n+2}$ and $I_{n+2}\times I_{n+1}$. We next show that the graphon is exactly constant on each rectangle. The proof is essentially a rerun of the analogous step in the proof of Theorem 6, only with more terms.
Let $g_{ij}$ be the average value of the graphon on the rectangle $I_{i}\times I_{j}$ and let $\Delta g_{ij}$ be the difference between the greatest and lowest value of the graphon in that rectangle. Let $c_{i}$ be the width of $I_{i}$. We have already determined that all $g_{ij}$’s except for $g_{n+1,n+2}$ and $g_{n+2,n+1}$ are exponentially close (in $\beta$) to 0 or 1, and hence that $1/H^{\prime\prime}(g_{ij})$ is exponentially small.
If $x\in I_{i}$ and $y\in I_{j}$, then the maximum and minimum possible values of $G(x,y)$, and their difference, are
(83) max $=$ $\sum_{k=1}^{n+2}c_{k}(g_{ik}+\Delta g_{ik})(g_{jk}+\Delta g_{jk}),$
(84) min $=$ $\sum_{k=1}^{n+2}c_{k}(g_{ik}-\Delta g_{ik})(g_{jk}-\Delta g_{jk}),$
(85) difference $=$ $\sum_{k=1}^{n+2}2c_{k}(g_{ik}\Delta g_{jk}+g_{jk}\Delta g_{ik}).$
Applying the mean value theorem to the Euler-Lagrange equations, and noting that $H^{\prime\prime}(u)$ is always negative, we have
(86) $-H^{\prime\prime}(g_{ij,0})\Delta g_{ij}\leq 2\beta\sum_{k=1}^{n+2}c_{k}(g_{ik}\Delta g_{jk}+g_{jk}\Delta g_{ik}),$
where $g_{ij,0}$ is some number between the values of $g$ corresponding to the maximum and minimum possible values of $G(x,y)$.
The sum on the right contains terms proportional to $\Delta g_{ij}$ itself, coming from $k=i$ or $k=j$. We bring those terms to the left hand side, noting that coefficients of those terms are much smaller than $H^{\prime\prime}(g_{ij,0})$. When $\{i,j\}\neq\{n+1,n+2\}$, this is because $g_{ij,0}$ is exponentially close to 0 or 1, so $H^{\prime\prime}(g_{ij,0})$ is exponentially large, while the coefficients on the right hand side are $O(\beta)$. When $\{i,j\}=\{n+1,n+2\}$, this is because the coefficients of $\Delta g_{n+1,n+2}$ on the right hand side are proportional to $\beta g_{n+1,n+1}$ or $\beta g_{n+2,n+2}$, both of which are exponentially small. By changing the factor of 2 on the right hand side to a 3, we can absorb these small corrections to the coefficient of $\Delta g_{ij}$ and also replace $H^{\prime\prime}(g_{ij,0})$ with just $H^{\prime\prime}(g_{ij})$. We also bound $g_{ik}$ and $g_{jk}$ by 1. The upshot is that
(87) $\Delta g_{ij}$$\leq$ $\frac{3\beta}{-H^{\prime\prime}(g_{ij})}\sum_{k}c_{k}(\Delta g_{jk}+\Delta g_{ik})$
(88) $\leq$ $\frac{6\beta}{-H^{\prime\prime}(g_{ij})}\max(\Delta g_{ik}\hbox{ or }\Delta g_{jk}),$
where the sum on the first line and the maximum on the second line skips terms involving $\Delta g_{ij}$ itself.
Whenever $\{i,j\}\neq\{n+1,n+2\}$, $\frac{6\beta}{-H^{\prime\prime}(g_{ij})}$ is exponentially small, so $\Delta g_{ij}$ is bounded by a tiny multiple of a sum of similar errors. In particular, the largest $\Delta g_{ij}$ of this sort is bounded by a sum of contributions much smaller than itself, possibly plus a contribution from $\Delta g_{n+1,n+2}$. The conclusion is that all $\Delta g_{ij}$’s other than $\Delta g_{n+1,n+2}$ are bounded by $\beta\exp(-\Omega(\beta))\Delta g_{n+1,n+2}$.
Now consider the equation for $\Delta g_{n+1,n+2}$. This equation indicates that $\Delta g_{n+1,n+2}$ is bounded by an $O(1)$ multiple of $\beta$ times the largest of the remaining $\Delta g_{ij}$’s, and so is bounded by a constant times $\beta^{2}\exp(-\Omega(\beta))\Delta g_{n+1,n+2}$. When $\beta$ is large, $\Delta_{n+1,n+2}$ is thus bounded by a constant (less than one) times itself, and so must be zero. But then all of the other $\Delta g_{ij}$’s must also be zero, so our graphon is multipodal.
4.5. Graphons with $(n,2)$ symmetry
Finally, we show that that the optimal graphon is symmetric in the first $n$ podes and symmetric in the last two. Let $c_{1},\ldots,c_{n+2}$ be the sizes of the various podes, let $\bar{c}$ be the average size of the first $n$ podes, and let $\Delta c_{i}$ be $c_{i}-\bar{c}$ or $c_{i}-\frac{1-n\bar{c}}{2}$, depending on whether we are talking about the first $n$ podes or the last two. Let $W_{i}$ be the worth of columns in the $i$-th pode. There are five kinds of rectangles, namely $I_{i}\times I_{j}$ with $i=j\leq n$, with $i<j\leq n$ or $j<i\leq n$, with $i\leq n<j$ or $j\leq n<i$, with $i=j>n$, and finally with $\{i,j\}=\{n+1,n+2\}$. In each class, let $\bar{g}_{ij}$ be the average value of the graphon and let $\Delta g_{ij}=g_{ij}-\bar{g}_{ij}$. We also refer to $g_{n+1,n+2}$ as $p$.
A key fact is that all of the entries in the first $n$ columns are exponentially close to 0 or 1. Meanwhile, the Euler-Lagrange equations for $g_{n+1,n+2}$ say that
(89) $H^{\prime}(p)$$\approx$ $\alpha+\beta\sum_{i=1}^{n}c_{i},$
(90) $\alpha$$\approx$ $H^{\prime}(p)-\beta\sum_{i=1}^{n}c_{i},$
where $\approx$” means “equal up to exponentially small corrections”.
Now suppose that $i$ and $j$ are indices less than or equal to $n$. Since all of the entries $g_{ik}$ and $g_{jk}$ are exponentially close to 0 or 1, the entropy contribution to $W_{i}$ or $W_{j}$ is exponentially small. The coefficient of $\alpha$ is $\sum_{k\neq i}c_{k}=1-c_{i}$, while the coefficient of $\beta/2$ is the integral of the graphon over everything that doesn’t involve the $i$-th pode. The upshot is that
(91) $W_{i}$$\approx$ $-\alpha(1-c_{i})-\frac{\beta}{2}(e-2c_{i}(1-c_{i}))$
(92) $=$ $-\left(\alpha+\frac{\beta e}{2}\right)+c_{i}(\alpha+\beta)-c_{i}^{2}\beta,$
with a similar result for $W_{j}$. Taking the difference gives
(93) 00 0 $=$ $W_{i}-W_{j}$
(94) $\approx$ $(c_{i}-c_{j})(\alpha+\beta-\beta(c_{i}+c_{j}))$
(95) $\approx$ $(c_{i}-c_{j})(H^{\prime}(p)+\beta(c_{n+1}+c_{n+2}-c_{i}-c_{j})),$
where we have used the fact that $\sum_{i=1}^{n}c_{i}=1-c_{n+1}-c_{n+2}$. However, $c_{n+1}$ and $c_{n+2}$ are close to $\frac{1-n\bar{c}}{2}$, while $c_{i}$ and $c_{j}$ are close to $\bar{c}$, so the coefficient of $\beta$ is bounded away from zero. We conclude that $c_{i}-c_{j}$ must be exponentially small. More precisely, $c_{i}-c_{j}$ must be exponentially smaller than the largest $|\Delta g_{ik}|$ or $|\Delta g_{jk}|$. A similar argument shows that $c_{n+1}-c_{n+2}$ is also exponentially smaller than the largest $|\Delta g|$.
We now look at the Euler-Lagrange equations for $g_{i\ell}$ and $g_{j\ell}$, where $\ell$ is different from $i$ or $j$. The difference between $G(x,y)$ in $I_{i}\times I_{\ell}$ and $I_{j}\times I_{\ell}$ is
(97) $\sum_{k=1}^{n+2}c_{k}g_{\ell k}(g_{ik}-g_{jk})$ $=$ $\sum_{k=1}^{n+2}c_{k}g_{\ell k}(\Delta g_{ik}-\Delta g_{jk})$
$+(c_{i}g_{\ell i}-c_{j}g_{\ell j})(\bar{g}_{ii}-\bar{g}_{ij}).$
The first line is of the order of the largest $\Delta g$. The second line has a similar contribution from the difference of $\Delta g_{\ell i}$ and $\Delta g_{\ell j}$, plus a contribution of order $c_{i}-c_{j}$. But then
(98) $H^{\prime}(g_{i\ell})-H^{\prime}(g_{j\ell})=\beta(G_{i\ell}-G_{j\ell}),$
which is $\beta$ times a linear combination of $\Delta g$’s and $c_{i}-c_{j}$. Since $H^{\prime\prime}$ is enormous on the interval from $g_{i\ell}$ to $g_{j\ell}$ (both of which are exponentially close to 1), $\Delta g_{i\ell}-\Delta g_{j\ell}$ is bounded by a tiny combination of other $\Delta g$’s and $\Delta c$’s.
Repeating this argument for $g_{ii}-g_{jj}$ and for $g_{n+1,n+1}-g_{n+2,n+2}$, we get that
  • The biggest $\Delta c_{i}$ is bounded by a tiny constant times the biggest $\Delta g$.
  • The biggest $\Delta g$ is bounded by a tiny constant times the biggest $\Delta c$.
We conclude that all of the $\Delta c$’s and $\Delta g$’s are zero.
We have determined the form of the optimal graphon $g$ at $(e,t)$. Noting that the Boltzmann entropy $\mathbb{B}(e,t)$ equals the Shannon entropy $S(g)$ of the optimal graphon $g$, all statements about $S$ are easily converted into statements about $\mathbb{B}$.
Finally, we must show that the values of $g$ on each rectangle, and the sizes of the different podes, are analytic functions of $(e,t)$. This follows from a general principle in algebraic geometry, which in turn is essentially just the implicit function theorem. Within the product of the Razborov triangle and the finite-dimensional space of graphons with $(n,2)$ symmetry, the set of optimal graphons is a 2-dimensional analytic variety, cut out by the analytic Euler-Lagrange equations. As long as the tangent space does not degenerate, we can write all but two of the variables as analytic functions of the last two, which we can choose to be $(e,t)$.
4.6. Distinct phases and rank
Theorem 8.
Each of the phases above the scallops proven in Theorems 6 and 7 have unique optimal graphons with distinct symmetries and cannot be analytically continued to one another.
Proof.
The optimal graphons described by Theorem 6 have rank 2, while the optimal graphons above the $n$-th scallop described by Theorem 7 have rank $n+2$. We will construct a sequence of “order parameters,” each a polynomial in finitely many subgraph densities, to distinguish between graphons of different rank. Specifically, the $k$th order parameter is identically zero whenever the rank of the optimal graphon is $k-1$ or less, and is never zero when the rank of the graphon is $k$. Since an analytic function on a connected set that is zero on an open subset is zero everywhere, there cannot be an analytic path connecting the $(k-2)$-nd scallop (where the graphon has rank $k$ and the order parameter is nonzero) to the previous scallops or to the $A(2,0)$ phase, where the order parameter is zero. In other words, the phases above the different scallops are all distinct.
Newton’s identities relate the determinant of a $k\times k$ matrix $A$ to the traces of $A^{j}$ for $j=1,2,\ldots,k$. For instance, if we let $t_{j}=\mathrm{Tr}(A^{j})$, then the determinants of small matrices are given by the formulas
(99) $\det(A)=\begin{cases}(t_{1}^{2}-t_{2})/2&k=2,\cr(t_{1}^{3}-3t_{1}t_{2}+2t_{3})/6&k=3,\cr(t_{1}^{4}-6t_{1}^{2}t_{2}+8t_{1}t_{3}-3t_{4})/24&k=4.\end{cases}$
Let $p_{k}(A)$ be the polynomial in the variables $\{t_{j}\}$ that gives the determinant of a $k\times k$ matrix $A$.
The same ideas work for arbitrary diagonalizable linear operators, for which the rank equals the number of nonzero eigenvalues, counted with multiplicity. If we evaluate $p_{k}$ on any diagonalizable trace-class operator, we get zero if the rank of the operator is less than $k$ and the product of the nonzero eigenvalues (counted with multiplicity) if the rank is equal to $k$. The key algebraic fact is that, for operators of rank $k$ or less, we have
(100) $t_{j}=\lambda_{1}^{j}+\cdots+\lambda_{k}^{j},$
where some of the eigenvalues $\lambda_{i}$’s may be zero, and $p_{k}$ computes $\lambda_{1}\cdots\lambda_{k}$, which is nonzero precisely when there are $k$ nonzero eigenvalues (counted with multiplicity).
In particular, we can apply these formulas to graphons. (Graphons are always diagonalizable, being symmetric and trace class.) For instance, the expression $(t_{1}^{4}-6t_{1}^{2}t_{2}+8t_{1}t_{3}-3t_{4})/24$, where now $t_{j}=\mathrm{Tr}(g^{j})$, gives zero if the rank of the graphon $g$ is less than 4 and gives a nonzero number if the rank is equal to 4.
When $j>2$, $t_{j}$ is the density of $j$-gons. The problem is that we cannot realize $t_{1}$ and $t_{2}$ as subgraph densities, so we cannot assume a priori that $t_{1}$ and $t_{2}$ are analytic functions of $(e,t)$ in each phase. To get around this problem, we define our $k$th order parameter to be $p_{k}(g^{3})$. This is still a polynomial in $\{t_{j}\}$, only now $j$ ranges from $3$ to $3k$ in steps of 3. In particular, $t_{j}$ is the density of $j$-gons for each applicable $j$.
The $k$-th order parameter is then zero if $g^{3}$ has rank less than $k$ and is nonzero if $g^{3}$ has rank $k$. But $g^{3}$ has the same rank as $g$, so we are actually testing the rank of $g$.
In summary: the $k$th order parameter is an analytic function of $(e,t)$ in each phase, being built from subgraph densities. It is identically zero on the regions above the 0th, 1st, $\ldots$, $(k-3)$rd scallops but is never zero on the region above the $(k-2)$nd scallop. Thus the region above the $(k-2)$nd scallop is in a different phase from the regions above all the previous scallops. Each scallop has its own unique phase.
5. Proof of Theorem 4
Once again we prove a slightly stronger version of the theorem stated in the introduction.
Theorem 9.
For each fixed $e\in(0,1)$ and all $t$ sufficiently close to (but below) $e^{3/2}$, the optimal graphon with edge/triangle densities $(e,t)$ is unique and bipodal. Asymptotically, the Boltzmann entropy scales as $-(e^{3/2}-t)\ln(e^{3/2}-t)$ and the Lagrange multiplier $\beta$ scales as $\ln(e^{3/2}-t)$.
Proof.
We follow the same overall roadmap as the proofs of Theorems 6 and 7. Specifically,
  1. (1)
    Using the proximity to the upper boundary, we break $[0,1]$ into two large podes $I_{1}$ and $I_{2}$ and a small exceptional set $I_{3}$ such that $g_{x}$ is $L^{2}$-close to the indicator function of $I_{1}$ when $x\in I_{1}$ and is $L^{2}$-close to zero when $x\in I_{2}$.
  2. (2)
    Equating the worths of $g_{x}$ when $x\in I_{1}$ to those of $g_{x}$ when $x\in I_{2}$, we determine that $\beta/\alpha\approx-2/e$. The multiplier $\beta$ is large and negative, while $\alpha$ is large and positive.
  3. (3)
    Maximizing $W(C)$ for an arbitrary $C:[0,1]\to[0,1]$, we show that every column is close to a typical column in the first or second pode. After reassigning points, $I_{3}$ is then empty. The control this gives us on $G(x,y)$ shows that $g(x,y)$ is everywhere exponentially close to 0 or 1.
  4. (4)
    Bounding the fluctuations in each rectangle by multiples of the fluctuations in other rectangles to show that all fluctuations are in fact zero. In other words, our optimal graphon is exactly bipodal with values that are exponentially close to 0 or 1.
Step 1 is identical to what we have done before. There is a unique graphon at $(e,e^{3/2})$, namely a graphon $g_{0}$ that is 1 on $I_{1}\times I_{1}$ and zero elsewhere, where $I_{1}$ is a pode of size $\sqrt{e}$. Every graphon with $t$ close to $e^{3/2}$, and in particular any entropy-maximizing graphon, must be $L^{2}$ close to $g_{0}$. This means that for all $x$’s outside of a set of small measure, $g_{x}$ is $L^{2}$-close to the corresponding column of $g_{0}$. This also implies that $G(x,y)$ is close to $\sqrt{e}$ when $x$ and $y$ are both in $I_{1}$ and is close to zero when either is in $I_{2}$.
The worth of a column that is nearly zero is of course nearly zero. The worth of a column that is nearly 1 on $I_{1}$ and nearly zero elsewhere is approximately
(101) $-\alpha\sqrt{e}-\frac{\beta e^{3/2}}{2}.$
Since all columns must have the same worth, we must have $\beta/\alpha\approx-2/e$. The Lagrange multipliers $\alpha$ and $\beta$ diverge at the same rate as we approach the boundary, with $\alpha\to\infty$ and $\beta\to-\infty$.
Now consider an arbitrary function $C:[0,1]\to[0,1]$. Let $a$ be the average of $C(y)$ on $I_{1}$ and let $b$ be the average on $I_{2}$. Using the approximation that $G$ is $L^{2}$-close to $\sqrt{e}$ times the indicator function of $I_{1}\times I_{1}$, we get that
(102) $W(C)\leq\sqrt{e}H(a)+(1-\sqrt{e})H(b)-\alpha\sqrt{e}a-\alpha(1-\sqrt{e})b-\frac{\beta}{2}e^{3/2}a^{2},$
with equality if $C$ is constant on $I_{1}$ and constant on $I_{2}$. Since $\alpha$ is large and positive, we must have $b$ exponentially close to 0. Setting $b\approx 0$, our worth is then approximately
(103) $\alpha\sqrt{e}(a^{2}-a).$
This is of course maximized at the endpoints $a=1$ and $a=0$, being negative when $a\in(0,1)$. In other words, any worth-maximizing column must either have $a\approx 1$ and $b\approx 0$, and so must be close to the columns in $I_{1}$, or $a\approx 0$ and $b\approx 0$, and so must be close to the columns in $I_{2}$. Reassigning the points of $I_{3}$ to $I_{1}$ or $I_{2}$ accordingly, we obtain a situation where $I_{3}$ is empty.
To constrain the fluctuations in $g(x,y)$ in each rectangle, we recall the variational equations
(104) $H^{\prime}(g(x,y))=\alpha+\beta G(x,y).$
Since $G(x,y)\approx 0$ or $\sqrt{e}$, depending on which quadrant we are in, this implies that $g(x,y)$ is exponentially close to 1 on $I_{1}\times I_{1}$ and exponentially close to 0 on $I_{1}\times I_{2}$ and $I_{2}\times I_{2}$. In particular, $H^{\prime\prime}(g)$, which scales as the larger of $1/g$ and $1/(1-g)$, is much larger than $|\beta|$. Looking at the change in the left hand side and right hand side of this equation within a single quadrant, we see that $H^{\prime\prime}(g)$ times the maximum fluctuation within any quadrant is of the same order as $|\beta|$ times the maximum fluctuation within any quadrant. But that means that the maximum fluctuation is bounded by a small multiple of itself, and so must be zero. Our graphon is exactly bipodal.
Finally, we do some calculations in the space of bipodal graphons. Let $g_{11}$, $g_{12}$ and $g_{22}$ be the values of the optimal graphon on $I_{1}\times I_{1}$, $I_{1}\times I_{2}$ and $I_{2}\times I_{2}$, respectively. We treat $g_{11}$, $g_{12}$ and $g_{22}$ as free variables and adjust the width of $I_{1}$ to keep the edge density fixed. To leading order, $e^{3/2}-t$ is a linear function of $(1-g_{11},g_{12},g_{22})$. However, $1-g_{11}$, $g_{12}$ and $g_{22}$ all scale as exponents of $\beta$, so $\beta$ must scale as $\ln(e^{3/2}-t)$. The entropy goes as $-(1-g_{11})\ln(1-g_{11})-g_{12}\ln(g_{12})-g_{22}\ln(g_{22})$, which then scales as $-(e^{3/2}-t))\ln(e^{3/2}-t)$.
The analyticity of $g$ as a function of $(e,t)$ follows from the same argument as in the proof of Theorem 7, only with the space of $(n,2)$ symmetric graphons replaced by the space of bipodal graphons.
6. Inequivalence of ensembles
Using Lagrange multipliers is superficially similar to studying an exponential random graph model (ERGM; see Section 6 in [10], or [14], or [12] for a relevant introduction), where one starts with values $\alpha$ and $\beta$ and defines a free energy
(105) $\Psi(\alpha,\beta)=\max_{g}\left(S(g)-\alpha\varepsilon(g)-\frac{\beta}{3}\tau(g)\right).$
(In the literature, ERGMs are usually described in terms of parameters $\beta_{1}=-\alpha$ and $\beta_{2}=-\beta/3$, but that linear rescaling is not important.) The function $\Psi(\alpha,\beta)$ is the Legendre transform of the Boltzmann entropy $\mathbb{B}(e,t)$:
(106) $\Psi(\alpha,\beta)=\max_{e,t}\left(\mathbb{B}(e,t)-\alpha e-\frac{\beta}{3}t\right).$
An LDP relates the graphon that maximizes the right hand side of (105), in other words that maximizes $F$ for the given values of $\alpha$ and $\beta$, to typical large graphs in a certain ensemble. In that alternate ensemble, all graphs on $n$ vertices are allowed, only with probability proportional to $\exp(-n^{2}(\alpha\varepsilon+\frac{\beta}{3}\tau))$, where $\varepsilon$ and $\tau$ are the edge and triangle densities of the graph in question.
If the Boltzmann entropy function were convex and the Razborov triangle were convex, then the Legendre transform (106) would be invertible within each phase [46]. If that were true, we could tune $\alpha$ and $\beta$ within each phase to get whatever values of $(e,t)$ we wanted, much as we can tune the temperature and pressure of a physical material to get whatever energy density and particle density that we want. In statistical mechanics, this ability to switch back and forth between fundamental variables and conjugate variables is called equivalence of ensembles. (Note that the equivalence of ensembles only involves invertibility in the interior of phases, not on the boundaries. The same temperature and pressure can describe a variety of water/ice mixtures with a range of energies.)
With graphs, the Boltzmann entropy function is not convex and neither is the Razborov triangle. There is no equivalence of ensembles; the Legendre transform (106) is not invertible. Specifically, there are many values of $(e,t)$ for which there do not exist any values of $(\alpha,\beta)$ whose $F$-maximizing graphons have edge/triangle densities $(e,t)$. We call such points $(e,t)$ ERGM-invisible. We can still understand $\Psi(\alpha,\beta)$ and the phases of an ERGM by studying $\mathbb{B}(e,t)$, since $\Psi$ is still the Legendre transform of $\mathbb{B}$. However, we cannot understand $\mathbb{B}(e,t)$, or graphs with general densities $(e,t)$, by studying $\Psi(\alpha,\beta)$.
It has long been known that all points with $t$ greater than $e^{3}$, and even moderately less than $e^{3}$, are ERGM-invisible [12]. The only points off the Erdős-Rényi curve that might be ERGM-visible lie close to the lower boundary of the triangle. We now show that, because of the nonconvexity of the Razborov triangle, most of those are also ERGM-invisible.
Theorem 10.
If $n$ is a positive integer and $\frac{n}{n+1}<e<\frac{n+1}{n+2}$, and if $t$ is sufficiently close to the minimum triangle density $t_{0}$, then $(e,t)$ is ERGM-invisible.
Proof.
Fix a value of $e$ strictly between $\frac{n}{n+1}$ and $\frac{n+1}{n+2}$. The points with edge density $e$ and triangle density just above the minimum must have large values of the Lagrange multipliers $\alpha$ and $\beta$. However, the scallop itself is concave down, so for positive values of $\beta$, the linear function $\alpha e+\beta t$ is greater at one or both of the neighboring cusps (at edge density $\frac{n}{n+1}$ and $\frac{n+1}{n+2}$) than near the interior of the scallop. For large positive values of $\beta$ (and correspondingly large negative values of $\alpha$), this difference is greater than the bounded difference in Shannon entropy between the graphons described by Theorem 7 and the zero-entropy graphons at the cusps. Since large values of $\beta$ correspond to small values of $\Delta t$, we conclude that, for all sufficiently small values of $\Delta t$, the point $(e,t_{0}+\Delta t)$ is ERGM-invisible. ∎
A similar result applies at the top of the Razborov triangle.
Theorem 11.
For each $e\in(0,1)$ and for all $t$ sufficiently close to (and less than) $e^{3/2}$, $(e,t)$ is ERGM-invisible.
Proof.
Near the top boundary, $\alpha$ is large and positive while $\beta$ is large and negative. However, the boundary curve $t=e^{3/2}$ is concave up, so we can decrease $\alpha e+\beta t/3$ by moving to one endpoint $(0,0)$ or the other $(1,1)$. Whenever $\alpha$ and $\beta$ are big enough in magnitude, in other words whenever we are close enough to the top boundary, these gains swamp any changes in $S(g)$. Either the constant graphon $g=0$ or the constant graphon $g=1$ yields a larger value of $F$ than the optimal graphon at $(e,t)$. ∎
The statement of Theorem 11 is not new; see Theorem 6.2 in [12]. However, the simplicity of the proof gets to the heart of why points in this region are ERGM-invisible.
Considering the Razborov triangle as a whole, only a few pieces are ERGM-visible. The Erdős-Rényi curve $t=e^{3}$ is ERGM-visible. Since $d\varepsilon$ and $d\tau$ are collinear at constant graphons, each point on the Erdős-Rényi curve actually corresponds to an infinite set of $(\alpha,\beta)$ values. A neighborhood of each cusp is ERGM-visible; that’s what you get when $\beta$ is large and positive. As far as we can tell, a neighborhood of the flat part of the lower boundary is also ERGM-visible. But that’s all. It appears that ERGMs can be used to study Erdős-Rényi graphs and parts of the $A(n,0)$ phases, but not the $F(1,1)$, $B(n,1)$ or $C(n,2)$ phases of Figure 2.
Acknowledgment. The work of L.S. was partially supported by the National Science Foundation under grant DMS-2113468.