### v3

# Emergence in Graphs with near-extreme constraints

Abstract.
We consider entropy-optimal graphons associated with extreme and near-extreme constraints on the densities of edges and triangles. We prove that the optimizers for near-extreme constraints are unique and multipodal and are perturbations of the previously known unique optimzers for extreme constraints. This proves the existence of infinitely many phases. We determine the podal structures in these phases and prove the existence of phase transitions between them.
1. Setting and results
In this paper we determine the structure of typical large graphs with a specified density $e$ of edges and a near-extreme density $t$ of triangles. It is known that the structure of all but an exponentially small fraction of such graphs are described by a graphon that maximizes a certain entropy functional subject to certain integral constraints, assuming that the entropy maximizer is unique up to equivalence. (See Section 1.3 for relevant background.) We show that this graphon is indeed unique and takes a simple multipodal form, with parameters that are piecewise analytic functions of $(e,t)$. We thereby prove the existence of infinitely many phases and infinitely many phase transitions. These results are stated in Theorems 1, 2, 3 and 4.
Our solutions to the entropy maximization problem rely on the theory of Lagrange multipliers and on a new quantity, called “worth”, that applies to columns of graphons. In Section 2 we develop the theory of Lagrange multipliers for graphons, define worth, and prove in Theorem 11 that all of the columns of an entropy-maximizing graphon must maximize worth.
1.1. The background
The boundary curves in Figure 1 show the extreme accessible values of pairs $(\varepsilon,\tau)$ of densities of edge and triangle subgraphs in asymptotically large simple graphs [45]. Values throughout the interior are also accessible, and can be studied, using the formalism of graph limits or graphons developed in 2006-2011 by Lovász et al [7, 8, 30, 31, 32] (see [4, 17] and [29] for background), and the Large Deviation Principle (LDP) of $\mathbb{G}(n,p)$ graphs [14].
Figure 1. The Razborov triangle, made from curves displaying the extreme values of pairs of accessible edge and triangle densities. The curvature of the “scallops” on the lower right is exaggerated for visibility.
The structure of the graphs on the boundary of Figure 1 appeared in 2012 in a preprint of [40]. Following as it did soon after the development of graphons and the LDP for $\mathbb{G}(n,p)$, this led to a series of papers aimed at analyzing emergent (large scale) structures when the edge/triangle constraints $(\varepsilon,\tau)$ were near the “Erdős-Rényi” curve $\tau=\varepsilon^{3}$, associated with graphs in $\mathbb{G}(n,p)$. These papers used the LDP to introduce a Boltzmann entropy, $\mathbb{B}(e,t)$, which measures the exponential rate of growth of the number of large graphs with edge and triangle densities $(\varepsilon,\tau)\approx(e,t)$, a notion copied from statistical mechanics. See [43] for a precise statement and proof of this theorem, and the important fact that $\mathbb{B}(e,t)$ is not convex in $(e,t)$.
Computer simulations led to the conjecture [26] of many distinct emergent “phases” throughout the interior of Figure 1, as seen in Figure 2. Eventually emergent phases were determined for parts of the three regions $F(1,1)$ [25], $B(1,1)$ [36] and $A(2,0)$ [37] in Figure 2 near the Erdős-Rényi curve.
Figure 2. Schematic drawing of a conjecture from 2017 [26], based on computer simulations of entropy-optimal graphons associated with the phases of large graphs with edge and triangle constraints.
1.2. Connections with other research
There is widespread interest in analyzing network and graph data [38], and statistical methods have developed to support this, including a family of Exponential Random Graph Models (ERGMs). In [13] a number of practical problems with ERGMs were isolated and treated using the large deviations principal of Erdős-Rényi graphs. In [42] a Boltzmann entropy $\mathbb{B}(\bar{\tau})$ was introduced which, together with the LDP, allowed the analysis of ‘exponentially most’ large finite graphs with constraints on the densities $\bar{\tau}$ of some subgraphs; see [43] for a review of these results in a general setting. Using $\mathbb{B}(\bar{\tau})$ some of the weaknesses of ERGMs discussed in [13] can be quantified. This is discussed in detail in Section 6.
Our paper is concerned with the emergence [28] of large scale structure in graphs, for instance of podes. By far the most highly developed mathematical formalism concerned with emergence is equilibrium statistical mechanics. In our work we have sought to create an analogous formalism for simple constrained graphs, motivated by the Razborov triangle, Figure 1.
One of our goals is to better understand constrained graphs, to extend the classic study of extremal graphs, of which [40] is a distinguished result [2, 6]. Another is to provide an example for other optimization problems facing similar obstacles. For this reason we give here a brief sketch of emergence in statistical mechanics [47, 53, 52], emphasizing the significance of entropy, free energies and especially convexity.
A common structure underlying the mathematics of edge-triangle graphs and the mathematics of statistical mechanics is constrained optimization. There are several ways to view equilibrium statistical mechanics as constrained optimization on a space of many-particle configurations [23, 19, 46, 27], all of which involve the global optimization of any of a range of free energy functionals, or the entropy. The entropy in statistical mechanics is a measure of the number of possible particle configurations with given constraints. It is a fundamental quantity. It is no exageration to view statistical mechanics as built on the convexity of this entropy (see the lectures of Lanford in [27] and the introduction by Wightman in [23]).
The convexity of the entropy allows one to analyze the system without loss of information by the use of a variety of free energies [52], such as the Gibbs free energy $G(p,T)$, which is more familiar than the entropy.
Figure 3. This is a crude sketch of the phases of bulk matter, separated by transition curves. There are more than 20 known different solid phases of water, different crystalline structures.
Figure 3 shows a primitive thermodynamic phase diagram, illustrating the pattern of solid and fluid phases in a physical bulk material (the phases represent the large particle-number scale in the emergent picture), as functions of constraint parameters pressure $p$ and temperature $T$ [41]. The experimentally measurable Gibbs free energy $G(p,T)$ [53, 48] is found to vary smoothly within each region, but shows singular behavior when constraints cross some lower dimension curves (see [46] and section VI in [23]), where bulk material properties such as mass density and heat capacity can change abruptly.
In the analogy between constrained graphs and statistical mechanics, discussed for instance in [39], [50] and [13], a large simple graph $G$ contains many edges, which play the role of individual particles, and the number of copies of some subgraph $H$ in $G$, such as triangles, plays the role of the total (potential) energy of $G$. A key to understanding emergent phases in both statistical mechanics and constrained graphs is defining an order parameter [3], which is a function that is identically zero in one phase but nonzero in another. We show its use in Section 4.6.
Unfortunately, with graphs there is no analog of equivalent free energies. (This is analyzed in Section 6.) This ‘inequivalence of ensembles’ presents a serious obstacle. In statistical mechanics free energies provide considerable technical advantages. In their absence we had to develop replacement tools. One such tool, which we introduce in this paper, is the “worth” $W(C)$ that we associate to each column $C$ of a graphon.
We note that some important optimization problems also suffer from nonconvexity complications. For instance, the form of optimal transport theory developed by Kantorovich and Brenier was based on convex analysis. It required significant developments over several years to allow the original nonconvex Monge problem to access that convex analysis. (For an introduction to optimal transport see the preface in [51] or chapter 1 in [33]).
The emergence of structure in large but finite physical materials is dramatic, particularly the diversity of solid phases, such as the graphite and diamond phases of carbon. (There are more than 20 distinct solid phases of water.) All the richness displayed by the phases of all materials, not just the pure elements but also the huge number of compounds like water and alcohol, is created by the electromagnetic interaction within and between molecules. All of inorganic chemistry comes down to the 100 or so different (integer) electric charges of atomic nuclei.
Emergence gives rise to the complicated structure of “water” in Figure 3 and in the edge-triangle model in Figure 2. The study of the emergence of such diversity from the interaction of invisibly small components of a small number of types has led to a great deal of interesting mathematics in the past twenty years. It is this promising history which was the motivation to bring the richness of emergent structure in statistical mechanics, which is built on the convexity of its entropy, into the nonconvex setting of the Boltzmann entropy $\mathbb{B}$ of constrained graphs.
1.3. Definitions
We include here the definitions and concepts needed to state our results precisely and to prepare the reader to follow the proofs.
1.3.1. Graphons and the cut topology
asdf
Intuitively, graphons are limits, as node number $n\to\infty$, of the adjacency matrices of simple graphs, with the nodes mapped to the interval $[0,1]$. For background up to its publication in 2011 and an encylopedic treatment of various aspects of graphons we recommend the book [29] by Lovász. For a more recent treatment, see [10].
Definition of graphons. The space $\mathcal{W}$ of graphons is the quotient, of the space of Borel measurable functions $g:[0,1]^{2}\to[0,1]$ satisfying $g(x,y)=g(y,x)$, by the identification of functions that differ only on a set of Lebesgue measure zero.
The cut topology on $\mathcal{W}$ is metrizable with the following metric.
Definition of the cut metric.
(1) $d_{cut}(g_{1},g_{2})\equiv\sup_{S,T}\Big{|}\int_{S\times T}\big{[}g_{1}(x,y)-g_{2}(x,y)\big{]}\,dx\,dy\Big{|},$
where $S,T$ are measurable subsets of $[0,1]$.
Note that the absolute value goes outside the integral!
1.3.2. Subgraph densities and weak equivalence
asdf
For a simple graph $F=(V,E)$ on $k$ nodes, we define the (homomorphism) density of $F$ in the graphon $g$ as
(2) $\tau_{F}(g)=\int_{[0,1]^{k}}\prod_{{i,j}\in E}g(x_{i},x_{j})dx_{1}\,dx_{2}\,\ldots,dx_{k}.$
It can be proven [10] that $\tau(F)$ is continuous on $\mathcal{W}$ in its cut topology. We will concentrate on the densities
(3) $\varepsilon(g)=\iint g(x,y)\,dx\,dy$
and
(4) $\tau(g)=\iiint g(x,y)g(y,z)g(z,x)\,dx\,dy\,dz$
of edges and triangles, respectively.
The space $\mathcal{W}$ of ordinary graphons is not compact in the cut topology. To obtain a compact space $\widetilde{\mathcal{W}}$ we take the quotient of $\mathcal{W}$ by a weak equivalence relation associated with subgraph densities.
Definitions of weak equivalence and reduced graphons. Graphons $g_{1}$ and $g_{2}$ are weakly equivalent if $\tau_{F}(g_{1})=\tau_{F}(g_{2})$ for every simple subgraph $F$. Elements of the quotient space $\widetilde{\mathcal{W}}$ are called reduced graphons.
For our purposes, it is useful to use a different description of weak equivalence. The group of measure-preserving transformations of $[0,1]$ acts naturally on $\mathcal{W}$. If $g$ is a graphon and $\sigma$ is such a transformation, we define, in terms of any representative modulo measure zero:
(5) $g^{\sigma}(x,y)=g(\sigma(x),\sigma(y)).$
We say that $g$ and $g^{\sigma}$ are group equivalent. Note that $\tau_{F}(g^{\sigma})=\tau_{F}(g)$, thanks to a simple change-of-variables in the integral (2), so group equivalence implies weak equivalence.
We next define a pseudometric $\delta_{cut}$ on $\mathcal{W}$ that measures how far two graphons are from being group equivalent.
Definition of $\delta_{cut}$.
(6) $\delta_{cut}(g_{1},g_{2})=\inf_{\sigma_{1},\sigma_{2}}d_{cut}(g_{1}^{\sigma_{1}},g_{2}^{\sigma_{2}}).$
It can be proven [10] that $\delta_{cut}(g_{1},g_{2})=0$ if and only if $g_{1}$ and $g_{2}$ are weakly equivalent. The pseudometric $\delta_{cut}$ on $\mathcal{W}$ then descends to a metric $\delta_{cut}$ on $\widetilde{\mathcal{W}}$: $\delta_{cut}([g_{1}],[g_{2}])=\delta_{cut}(g_{1},g_{2})$, where $[g]$ denotes the weak equivalence class of $g\in\mathcal{W}$.
1.3.3. Compactness, the LDP of $\mathbb{G}(n,p)$ graphs and the Shannon entropy
asdf
A major result in the graphon formalism is the
Compactness theorem. $\widetilde{\mathcal{W}}$ is compact in the topology defined by the metric $\delta_{cut}$.
Our results rely heavily on the LDP of $\mathbb{G}(n,p)$ graphs, which is expressed in terms of graphons. For an elegant and concise reference for both the graphon formalism and the LDP we recommend the book [10] of Chatterjee. [^1] Note, however, that Chatterjee works with a slightly different metric on $\mathcal{W}$, defining $d^{\prime}_{cut}(g_{1},g_{2})=\sup_{a,b}\Big{|}\int_{[0,1]^{2}}a(x)b(y)\big{[}g_{1}(x,y)-g_{2}(x,y)\big{]}\,dx\,dy\Big{|},$ where $a$ and $b$ are Borel measurable maps from $[0,1]$ to [1,1][-1,1]. $d^{\prime}_{cut}(g_{1},g_{2})$ is bounded above and below by multiples of $d_{cut}(g_{1},g_{2})$, so Chatterjee’s topological results based on $d^{\prime}_{cut}$ and the corresponding $\delta^{\prime}_{cut}$ apply equally well to $d_{cut}$ and $\delta_{cut}$.
We next introduce some terms associated with the LDP.
The Shannon entropy of a graphon $g$ is
(7) $S(g)=\iint H\big{(}g(x,y)\big{)}\,dx\,dy,$
where
(8) $H(u)=-[u\ln(u)+(1-u)\ln(1-u)]$
is the usual entropy of independent coin flips with probability $u$ of getting heads. Note that $H(u)$ is concave down and that $H^{\prime}(u)=\ln\left(\frac{1-u}{u}\right)$ diverges as $u$ approaches 0 or 1. $S(g)$ is invariant under measure-preserving transformations of $[0,1]$ and defines a function (also denoted $S$) on $\widetilde{\mathcal{W}}$. $S$ is also minus the rate function of the LDP of $\mathbb{G}(n,p)$ graphs. The LDP relates the number of large graphs associated with an (open or closed) subset of $\widetilde{\mathcal{W}}$ with the supremum of $S$ on that subset; see [10] for more details on the LDP.
Let $\mathcal{W}_{e,t}$ be the set of graphons with $\varepsilon(g)=e$ and $\tau(g)=t$ and let $\widetilde{\mathcal{W}}_{e,t}$ be the corresponding set of reduced graphons.
The Boltzmann entropy function $\mathbb{B}(e,t)$ can be understood in two ways. One is as the exponential rate of growth, as the number of nodes $n$ diverges, of the number of graphs on $n$ nodes with edge/triangle densities $(e,t)$. For this paper, it is useful to use the fact, proven in [42, 44], that $\mathbb{B}(e,t)$ equals the maximum of $S([g])$ on $\widetilde{\mathcal{W}}_{e,t}$, which is the same as the maximum of $S(g)$ on $\mathcal{W}_{e,t}$.
Most of this paper is devoted to maximizing $S$ on $\mathcal{W}_{e,t}$ or $\widetilde{\mathcal{W}}_{e,t}$. That is, we focus on the constrained optimization of $S$. The optimizing graphons are called entropy-optimal graphons (or “optimal graphons”, for short) for the given constraints $(e,t)$. When we refer to an optimal graphon being unique we always mean that the optimal reduced graphon is unique.
The conjecture [44, 26] that constrained optimization of $S$ would give rise to the rich phenomena of Figure 2, much of which is finally proven in this paper, was by analogy of $\mathbb{B}(e,t)$ with the Boltzmann entropy in statistical mechanics, as discussed in subsection 1.2.
1.3.4. Multipodality, phases and phase transitions
asdf
A graphon is said to be k-podal if we can partition $[0,1]$ into $k$ measurable sets $I_{1},\ldots,I_{k}$ such that $g(x,y)$ is constant on each “rectangle” $I_{i}\times I_{j}$. We refer to the sets $I_{i}$ as podes. If $g$ is $k$-podal for some integer $k$, we say that $g$ is multipodal. We often use the words bipodal and tripodal to mean 2-podal and 3-podal. We say that an $(n+m)$-podal graphon has $(n,m)$ symmetry if $g$ is invariant under permutation of $n$ of the podes and is also invariant under permutation of the remaining $m$ podes. A graphon with $(2,0)$ symmetry is said to be symmetric bipodal.
A phase is a connected open set in the interior of the Razborov triangle (Figure 1) in which the reduced optimal graphon is unique and is a real analytic function of $(e,t)$ in the following sense. Within a phase, and for each simple graph $F$, $\tau_{F}$ of the optimal graphon is required to be an analytic function of $(e,t)$. In practice, the analyticity of all densities $\tau_{F}$ is proven by first showing that the optimal graphon is unique and multipodal of a certain form and then showing that the finitely many parameters needed to describe this multipodal graphon are analytic functions of $(e,t)$.
A phase transition occurs where some $\tau_{F}$ is not analytic or is not even defined, such as where the optimal graphon is not unique. Phase transitions have only been shown to occur on boundaries of phases, on curves.
1.4. Detailed results
The following theorems give a simple description of what happens near almost all points along the boundary of the Razborov triangle, Figures 1 and 2. We prove that the unique optimal graphons in the phases near the boundary are multipodal. The cited theorems in later sections include additional estimates on how the parameters of the optimal graphons scale as the constraints approach the boundary.
Theorem 1 (Theorem 14).
For each fixed $e<1/2$ and all $t$ sufficiently small, the optimal (reduced) graphon in $\widetilde{\mathcal{W}}_{e,t}$ is unique and is symmetric bipodal, with parameters that vary analytically with $(e,t)$.
Theorem 2 (Theorem 17).
Let $n\geq 1$ be an integer. For every $e\in\left(\frac{n}{n+1},\frac{n+1}{n+2}\right)$, with corresponding minimal triangle density $t_{0}$ (depending on $e$), and for all $\Delta t$ sufficiently small, the optimal (reduced) graphon in $\widetilde{\mathcal{W}}_{e,t_{0}+\Delta t}$ is unique and $n+2$-podal, with $(n,2)$ symmetry and with parameters that vary analytically with $(e,t)$.
Note that these theorems do not make any claims about what happens exactly over the cusps, i.e. when $e=\frac{n}{n+1}$. When $e=1/2$ ($n=1$), the optimal graphon has long been known to have a symmetric bipodal structure. When $n$ is larger, the optimal graphon is believed to have $(n+1,0)$ symmetry. However, a small neighborhood of each cusp is believed to intersect four(!) different phases, making a precise characterization difficult.
Theorem 3 (Theorem 19).
All of the phases above the scallops proven in Theorems 1 and 2 have unique optimal reduced graphons with distinct symmetries and cannot be analytically continued to one another.
In the notation of Figure 2, Theorem 1 proves that the region just above the flat part of the bottom boundary is part of the $A(2,0)$ phase. Theorem 2 proves the existence of all of the $C(n,2)$ phases, and Theorem 3 shows that these phases are all different.
Theorem 4 (Theorem 20).
For each fixed $e\in(0,1)$ and all $t$ sufficiently close to (but below) $e^{3/2}$, the optimal graphon with edge/triangle densities $(e,t)$ is unique and bipodal, with parameters that vary analytically with $(e,t)$.
That is, the region just below the upper boundary is part of a bipodal phase. There is every reason to believe that this is part of the same bipodal $F(1,1)$ phase that is found just above the ER curve, but this has not yet been proved.
All of these theorems can be viewed as extensions of extremal graph theory. Pikhurko and Razborov’s results [40] determine unique, entropy-optimal graphons on the boundary of Figure 1 [42]. Theorems 14 describe the infinite number of distinct neighboring phases.
2. Lagrange multipliers and the “worth” functional
In this section we develop the concept of “worth”. This is a quantity associated with columns of a graphon, that is with the functions $g_{x}(y)=g(x,y)$ of one variable defined by fixing $x$ and allowing $y$ to float. We will see in Theorem 11 that the columns of an entropy optimal graphon must maximize worth. Working with columns, rather than just with the value $g(x,y)$ of the graphon $g$ at each point $(x,y)$ separately, gives us the analytical control needed to prove our main theorems.
To accomplish this we need a theory of Lagrange multipliers. For these purposes, we treat graphons as elements of $L^{2}([0,1]^{2})$ with the $||\cdot||_{2}$-norm. The $L^{2}$-topology is finer than the topology defined by the cut distance. All $L^{2}$-limits are limits in the cut metric, but not vice-versa. In order to combine our machinery with established theorems about reduced graphons, we will eventually need to prove that certain sequences converge in $L^{2}$ and not just in the cut distance.
We develop Lagrange multipliers in several steps. First we develop the theory for a certain class of variations where the standard theory of functional derivatives applies. This determines our Lagrange multipliers $(\alpha,\beta)$. We then show that, for other $L^{2}$-small changes to an optimal graphon, a certain functional involving $\alpha$ and $\beta$ cannot increase to leading order in the size of the change. We compute the effect of changing a small set of columns of our graphon and express the difference in this functional in terms of the worths of the old and new columns. Since the functional cannot increase, we conclude that the columns of the original (optimal) graphon must all maximize worth.
The simplest changes involve varying the value of our graphon $g$ gradually at each point, as indicated by a bounded symmetric function $g_{1}:[0,1]^{2}\to\mathbb{R}$. We consider a family of graphons $g_{s}$ of the form
(9) $g_{s}(x,y)=g_{0}(x,y)+sg_{1}(x,y).$
To remain in $\mathcal{W}$ we must have $0\leq g_{0}(x,y)+sg_{1}(x,y)\leq 1$ for all sufficiently small $s$. This can always be achieved by choosing $g_{1}$ to be supported on a subset of $0,1]^{2}$ on which $g_{0}(x,y)$ is bounded away from 0 and 1. For now, we do not consider variations that change $g(x,y)$ at points where $g_{0}(x,y)=0$ or 1.
The resulting changes to the edge density, triangle density and Shannon entropy are:
(10) $\Delta\varepsilon:=\varepsilon(g_{s})-\varepsilon(g_{0})$ $=$ $s\iint g_{1}(x,y)\,dx\,dy+o(s),$
(11) $\Delta\tau:=\tau(g_{s})-\tau(g_{0})$ $=$ $s\iint 3G(x,y)g_{1}(x,y)\,dx\,dy+o(s),$
(12) $\Delta S:=S(g_{s})-S(g_{0})$ $=$ $s\iint H^{\prime}(g_{0}(x,y))g_{1}(x,y)\,dx\,dy+o(s),$
where
(13) $G(x,y)=\int_{0}^{1}g_{0}(x,z)g_{0}(y,z)\,dz,$
and $H^{\prime}(u)=\ln\left(\frac{1-u}{u}\right)$ is the derivative of $H$ defined in equation (8). The quantities $H^{\prime}(g_{0}(x,y))$, 11 and $3G(x,y)$ are called the functional derivatives of $S(g)$, $\varepsilon(g)$, and $\tau(g)$ with respect to $g$ at $g=g_{0}$ and are sometimes denoted
(14) $\frac{\delta S(g)}{\delta g(x,y)},\qquad\frac{\delta\varepsilon(g)}{\delta g(x,y)},\qquad\hbox{and}\qquad\frac{\delta\tau(g)}{\delta g(x,y)}.$
These functional derivatives are elements of $L^{2}(A)$, where $A$ is the subset of $0,1]^{2}$ where $0<g(x,y)<1$, and are only defined up to sets of measure zero.
Remark 5.
The graphon $g$ is an element of $L^{2}([0,1]^{2})$ and so is not literally a function. Rather, it is an equivalence class of functions that agree off sets of measure zero. Similarly, the “function” $G(x,y)$, which gives the inner product of the columns $g_{x}$ and $g_{y}$, is only defined up to sets of measure zero. If $\bar{g}:[0,1]^{2}\to[0,1]$ is a symmetric function that agrees with $g$ apart from a negligible set, then Fubini’s Theorem says that, for almost every $x$, the columns $g_{x}$ and $\bar{g}_{x}$ agree except on a negligible subset of $[0,1]$, and in particular represent the same “function” in $L^{2}([0,1])$. So for almost every pair $(x,y)\in[0,1]^{2}$, the inner product of $g_{x}$ and $g_{y}$ is the same as the inner product of $\bar{g}_{x}$ and $\bar{g}_{y}$. That is, the functions $G(x,y)$ computed from $g$ and $\bar{g}$ agree except on a set of measure zero.
Lemma 6.
If $g$ is not a constant graphon, then $G(x,y)$ is not constant.
Proof.
By Cauchy-Schwarz,
(15) $G(x,y)\leq\sqrt{G(x,x)G(y,y)}\leq\max(G(x,x),G(y,y)).$
Proving by contradiction, the only way for $G(x,y)$, $G(x,x)$ and $G(y,y)$ to all be equal is if the columns of $g$ at $x$ and $y$ are identical. But if all the columns of $g$ are identical (and likewise all of the rows, since $g$ is symmetric), then $g$ is a constant graphon. ∎
Theorem 7.
Let $g$ be an entropy-maximizing graphon, subject to the constraints $\varepsilon(g)=e$ and $\tau(g)=t$. If the function $G(x,y)$ is not constant on the set of points $(x,y)$ where $0<g(x,y)<1$, then there exist unique Lagrange multipliers $\alpha$ and $\beta$ such that
(16) $H^{\prime}(g(x,y))=\alpha+\beta G(x,y)$
almost everywhere. Furthermore, the function $g(x,y)$ is bounded away from 0 and 1.
Proof.
Since $G(x,y)$ is not constant on the set $0<g(x,y)<1$, $G(x,y)$ is not constant on the set $\epsilon<g(x,y)<1-\epsilon$ for all sufficiently small $\epsilon$, so the functional derivatives of $\varepsilon$ and $\tau$ are linearly independent on this set, so we can vary $\varepsilon$ and $\tau$ independently by choosing appropriate functions $g_{1}(x,y)$ that are supported on this set. Similarly, if the three functions $\frac{\delta\varepsilon}{\delta g}$, $\frac{\delta\tau(g)}{\delta g(x,y)}$, and $\frac{\delta S(g)}{\delta g(x,y)}$ were linearly independent on the set of points where $0<g(x,y)<1$, then we could vary $\varepsilon$, $\tau$, and $S$ independently, and in particular we could increase $S$ while keeping $e$ and $t$ exactly constant. Since this contradicts the optimality of $g$, we conclude that $\frac{\delta S(g)}{\delta g(x,y)}$ is a linear combination of $\frac{\delta\varepsilon}{\delta g}$ and $\frac{\delta\tau(g)}{\delta g(x,y)}$, so we can write
(17) $\frac{\delta S}{\delta g(x,y)}=\alpha\frac{\delta\varepsilon}{\delta g(x,y)}+\frac{\beta}{3}\frac{\delta\tau}{\delta g(x,y)}.$
Since 1 and $G(x,y)$ are linearly independent, the coefficients $\alpha$ and $\beta$ are uniquely defined. Plugging the functional derivatives into equation (17) then gives (16), which applies at points where $0<g(x,y)<1$.
Since $\alpha$ and $\beta$ are finite and since $G(x,y)$ is bounded, $H^{\prime}(g(x,y))$ is bounded, so $g(x,y)$ cannot be arbitrarily close to 0 or 1. Either $g(x,y)$ equals 0 or 1 (in which case equation (16) does not apply) or $g(x,y)$ is bounded away from 0 or 1.
We now eliminate the first possibility. If $g(x,y)=0$ or 1 on a set of positive measure, we can change the value of $g(x,y)$ at such points away from 0 and 1, thereby increasing $S$ greatly while only changing $\varepsilon(g)$ and $\tau(g)$ slightly. Since we can vary $\varepsilon$ and $\tau$ independently with an appropriate choice of $g_{1}$, we can restore the original values of $\varepsilon$ and $\tau$ with the resulting change in $S$ being governed by (17). The total effect is to increase $S$ while leaving $\varepsilon$ and $\tau$ fixed, which is a contradiction. ∎
The crux of the proof is the same as the derivation of Lagrange multipliers in finite dimensions. The fact that linearly independent derivatives give us the ability to vary $\varepsilon$ and $\tau$ (and possibly $S$) independently is just the inverse function theorem in $\mathbb{R}^{2}$ and $\mathbb{R}^{3}$. The treatment of the points where $g(x,y)=0$ or 1 is essentially the same as what we do to maximize a function on the boundary of a domain in $\mathbb{R}^{n}$. Our setting is the infinite-dimensional space of graphons, but the core arguments are just finite-dimensional calculus.
Theorem 7 was stated in terms of functional derivatives, but we can also speak in terms of small changes to a graphon. As long as the functional derivatives of $\varepsilon$ and $\tau$ are linearly independent on the set of points where $0<g(x,y)<1$, we can vary $\varepsilon$ and $\tau$ independently by choosing an appropriate $g_{1}(x,y)$, with the resulting change in $S$ being given by
(18) $\Delta S=\alpha\Delta\varepsilon+\frac{\beta}{3}\Delta\tau+o(s\|g_{1}\|_{2}).$
Remark 8.
But what if a graphon $g$ fails to meet the assumptions of Theorem 7, with $g(x,y)$ equaling 0 or 1 on part of the unit square and $G(x,y)$ being constant everywhere else? That is, what if we cannot vary $\varepsilon$ and $\tau$ independently at $g$? We call such graphons singular entropy maximizers. These certainly appear on the boundary of the Razbarov triangle. In the interior they appear on the Erdős-Rényi curve $t=e^{3}$ and may appear elsewhere, so sometimes extra work is needed to exclude them. (See [26, Theorem 4.1] for such an argument at a specific phase transition.)
Fortunately for us, such singular graphons do not appear in any of the regions being studied in this paper. We postpone an explanation of this fact to the end of this section and return to the more typical situation of nonsingular entropy maximizers that satisfy the assumptions of Theorem 7.
We now consider infinitesimal changes to a graphon obtained by changing $g(x,y)$ by a “macroscopic amount” on a set of infinitesimal measure $s$. For instance, we might vary the sizes of the podes in a multipodal graphon. Such changes are not covered by Theorem 7. Nonetheless, they satisfy Lagrange-like inequalities.
Proposition 9.
Suppose that $g_{0}$ is a non-constant and non-singular entropy maximizer with Lagrange multipliers $\alpha$ and $\beta$. Let $g_{s}$ be a family of graphons obtained by changing $g_{0}$ on sets of measure $s$. Then
(19) $\Delta S\leq\alpha\Delta\varepsilon+\frac{\beta}{3}\Delta\tau+o(s),$
where $\Delta S=S(g_{s})-S(g_{0})$, $\Delta\varepsilon=\varepsilon(g_{s})-\varepsilon(g_{0})$, and $\Delta\tau=\tau(g_{s})-\tau(g_{0})$.
Proof.
Supposing this to be false, we will construct a variation of $g_{0}$ that has the same edge and triangle densities but more entropy. As previously noted, we can adjust $\varepsilon$ and $\tau$ independently by adding a function supported on $\epsilon<g(x,y)<1-\epsilon$, with the resulting changes in $S$ given by equation (18). Applying these changes to $g_{s}$ to restore the original values of $\varepsilon$ and $\tau$, we should get a change in $S$ that is at least $\Theta(s)-o(s)>0$, which is a contradiction.
To complete this argument, we must bound the cross terms from adding a function of pointwise size $O(s)$ to $g_{0}$ and in changing $g_{0}$ on a set of measure $s$. Changing $g(x,y)$ by $O(s)$ can only change $H(g(x,y))$ by $O(s\ln(s))$ and can only change $g(x,y)g(y,z)g(x,z)$ by $O(s)$. Integrating over a region of size $O(s)$, this can change the entropy by $O(s^{2}\ln(s))$, the edge density by $O(s^{2})$ and the triangle density by $O(s^{2})$, resulting in an $o(s)$ change to $S-\alpha\varepsilon-\frac{\beta}{3}\tau$, which can be absorbed into the $o(s)$ error term in equation (18). ∎
The estimate (19), combined with the theory of Lagrange multipliers for pointwise-small changes (Theorem 7 as summarized in equation (18)) can be described in terms of a functional
(20) $F(g)=S(g)-\alpha\varepsilon(g)-\frac{\beta}{3}\tau(g).$
If $g$ is a nonsingular constrained entropy maximizer, then pointwise small changes to $g$ cannot change $F$ to first order, while macroscopic changes on small sets can decrease $F$ but cannot increase $F$ to first order.
The functionals $S$ and $\varepsilon$ are local, in that there is a contribution from each point $(x,y)$ in $0,1]^{2}$ and we integrate the local contributions to get the global quantity. If the triangle density were also local, then $F$ would be the integral of a local density and our variational equations would come from setting the derivative of this density with respect to $g(x,y)$ equal to zero. That is the typical situation when doing calculus of variations, especially in classical and quantum field theory, with the global action being the integral of a local Lagrangian density [5].
However, the triangle density
(21) $\tau(g)=\iiint g(x,y)g(y,z)g(z,x)\,dx\,dy\,dz$
is not local. It involves interactions between the values of $g$ at the three points $(x,y)$, $(x,z)$ and $(y,z)$. To accommodate this complication it is useful to consider macroscopic changes to entire columns. We define a quantity that measures the effect of such changes.
Definition of worth: Let $C$ be a possible column of a graphon $g$. That is, $C:[0,1]\to[0,1]$ is a Borel measurable function. The worth of $C$ is
(22) $W(C)=2\int_{0}^{1}H(C(y))\,dy-2\alpha\int_{0}^{1}C(y)\,dy-\beta\iint C(y)C(z)g(y,z)\,dy\,dz.$
Note that this depends explicitly on the graphon $g$ as well as on $C$ and the Lagrange multipliers $\alpha$ and $\beta$.
Proposition 10.
Let $(\alpha,\beta)$ be specified and suppose that $\tilde{g}(x,y)=g(x,y)$ except when $x$ or $y$ lies in a set $I$ of measure $s$. Then
(23) $F(g)-F(\tilde{g})=\int_{I}W(g_{x})-W(\tilde{g}_{x})\,dx+O(s^{2}),$
where $g_{x}$ and $\tilde{g}_{x}$ are columns of $g$ and $\tilde{g}$.
Proof.
The quantities $\varepsilon(g)$, $S(g)$ and $\tau(g)$ are all double or triple integrals over $0,1]^{2}$ or $0,1]^{3}$. The integrands for $g$ and $\tilde{g}$ are identical except where one of the variables lies in $I$. To compute $F(g)-F(g^{\prime})$, we must only keep the contributions of $x\in I$, multiply by 2 or 3 to allow for the similar contributions of $y\in I$ or $z\in I$, and make adjustments for where two or three variables are in $I$. Since the integrand is bounded and the set of points where multiple variables lie in $I$ only has measure $O(s^{2})$, we can compute F(g)F(g~)F(g)-F($\tilde${g}) to within $O(s^{2})$ by assuming that $x\in I$ and leaving $y$ and $z$ free.
We begin with the edge density:
(24) $\varepsilon(g)-\varepsilon(g^{\prime})=2\int_{x\in I}\int_{y\in[0,1]}g_{x}(y)-\tilde{g}_{x}(y)\,dy\,dx+O(s^{2}).$
The entropy is similar:
(25) $S(g)-S(\tilde{g})=2\int_{x\in I}\int_{y\in[0,1]}H(g_{x}(y))-H(\tilde{g}_{x}(y))\,dy\,dx+O(s^{2}).$
The triangle density has a prefactor of 3 instead of 2 because it is a triple integral:
[…]
(27) $=$ $3\int_{x\in I}\iint_{y,z\in[0,1]}\left(g_{x}(y)g_{x}(z)-\tilde{g}_{x}(y)\tilde{g}_{x}(z)\right)g(y,z)\,dy\,dz\,dx+O(s^{2}),$
where in the last line we used the fact that $\tilde{g}(y,z)$ only differs from $g(y,z)$ when $y\in I$ or $z\in I$. Multiplying the change in $\varepsilon$ by $\alpha$ and the change in $\tau$ by $\beta/3$ and adding terms, we get (23). ∎
Theorem 11.
If $g$ is a nonsingular entropy maximizer with Lagrange multipliers $(\alpha,\beta)$ then $g$ agrees off a set of measure zero with a graphon where every column maximizes $W$. In particular, every column must have the same worth.
Proof.
First we show that almost every column maximizes worth. Since we are working in $L^{2}$, where functions that differ on sets of measure zero are considered equivalent, we can get an equivalent representative by replacing any columns that don’t maximize worth with ones that do.
Let $W_{max}$ be the supremum of the worths of all possible columns. If there is a set of columns of positive measure whose worths are strictly less than $W_{max}$ then for some $\delta>0$ there is a set of columns of positive measure whose worths are all bounded above by $W_{max}-2\delta$. Replacing a subset $I_{s}$ of measure $s$ of such columns with a column $C$ whose worth is within $\delta$ of $W_{max}$ will increase $F$ by at least $s\delta+O(s^{2})$, contradicting Proposition 9.
(Note that “change a set of columns to CC” leads to an ambiguity for $g(x,y)$ when both $x$ and $y$ are in $I_{s}$. We can resolve this ambiguity by setting $g(x,y)=1$ on $I_{s}\times I_{s}$, or by setting $g(x,y)=0$, or by picking any other symmetric function in this square. Since the square where the ambiguity occurs has area $s^{2}$, different choices will yield values of $F$ that differ by $O(s^{2})$, which does not affect the violation of Proposition 9.) ∎
Having established the variational equations for nonsingular entropy maximizers, both for points with Theorem 7 and for columns with Theorem 11, we return to the (non)existence of singular entropy maximizers. If $g$ is a singular entropy maximizer, we call $\tau(g)$ a singular value of $t$ for the given edge density $e=\varepsilon(g)$. We claim that singular $t$’s are too sparse to matter. We begin with their measure.
Lemma 12.
For each $e\in(0,1)$, the set of singular $t$-values has measure zero.
Proof.
As explained below, the Boltzmann entropy function $\mathbb{B}(e,t)=\max_{g\in\mathcal{W}_{e,t}}S(g)$ for fixed $e$ is monotonically increasing in $t$ on $(t_{min},e^{3})$ and monotonically decreasing on $(e^{3},e^{3/2})$. This makes $d\mathbb{B}$ a finite measure on $(t_{min},e^{3})$ and makes $d\mathbb{B}$ a finite measure on $(e^{3},e^{3/2})$. We will show that the measure is singular at all singular $t$-values. The theorem then follows from the fact that the support of the singular part of a finite measure on an interval has zero Lebesgue measure.
To see the monotonicity of $\mathbb{B}$, let $g_{0}$ be an entropy maximizer at $(e,t)$ with $t<e^{3}$ and consider the family of graphons
(28) $g_{s}(x,y)=se+(1-s)g_{0}(x,y),$
which is defined for all $s\in[0,1]$. $S(g_{s})$ is an increasing function of $s$ (thanks to the concavity of $H(u)$) and in particular $S(g_{s})>S(g_{0})$ for all $s>0$. Since $\tau(g_{s})$ goes from $t$ to $e^{3}$ as $s$ goes from 0 to 1, and since $S(g_{s})$ is a lower bound for $\mathbb{B}(e,\tau(g_{s}))$, $\mathbb{B}(e,t^{\prime})>\mathbb{B}(e,t)$ for every $t^{\prime}\in(t,e^{3})$.
We also note that $\tau(g_{s})>t$ for all $s>0$, since otherwise, by the intermediate value theorem, there would be a positive $s$ with $\tau(g_{s})=t$. Since $S(g_{s})>S(g_{0})$, as we just discussed, that would contradict the optimality of $g_{0}$.
Now suppose that $t$ is singular, with a singular entropy maximizer $g_{0}$. Since $g_{0}(x,y)$ equals 0 or 1 on a set of positive measure, $S(g_{s})-S(g_{0})$ scales as $s\ln(1/s)$ as $s\to 0$. However, $\tau(g_{s})$ is a polynomial in $s$ and cannot grow faster than linearly for small $s$. Thus
(29) $\lim_{s\to 0^{+}}\frac{S(g_{s})-S(g_{0})}{\tau(g_{s})-\tau(g_{0})}=+\infty.$
Since $\mathbb{B}(e,t)=S(g_{0})$, and since $S(g_{s})$ is a lower bound for $\mathbb{B}(e,\tau(g_{s}))$, $\mathbb{B}$ must be increasing infinitely fast at $t$. That is, $t$ is a singular point of the measure $d\mathbb{B}$.
The exact same arguments work for $t>e^{3}$, only with $\tau(g_{s})$ being a decreasing function of $s$, with $\mathbb{B}$ being a decreasing function of $t$, and with $\mathbb{B}$ decreasing at infinite rate at singular $t$-values. ∎
We note that Lemma 12 proves there are no singular entropy maximizers in the regions studied in this paper.
Theorem 13.
Fix the edge density $e$ and consider an open interval $I$ of triangle densities $t$. Suppose that there is a differentiable function $f(t)$, defined for all $t\in I$, that equals the Boltzmann entropy $\mathbb{B}(e,t)$ whenever $t$ is non-singular. Then every $t\in I$ is non-singular.
Proof.
The Boltzmann entropy is never differentiable at $t=e^{3}$ [44], so we only need to consider intervals $I$ that are either above or below the Erdős-Rényi curve. By Lemma 12, the set of singular $t$-values in $I$ has measure zero, so the complement of that set is dense in $I$. Since $\mathbb{B}(e,t)$ is monotonic on $I$ and equals a (differentiable and therefore) continuous function $f(t)$ on a dense subset of $I$, it must equal $f(t)$ on all of $I$. But then $\mathbb{B}(e,t)$ is differentiable in $t$ for all $t\in I$, so there are no singular $t$-values in $I$.
Theorems 11 and 13 give us a strategy by which we can now prove (with much work!) Theorems 1, 2 and 4. In each case, we use variational equations that apply whenever $t$ is non-singular. For all such $t$, we show that the optimal graphon must take a certain form, with an entropy that is (the restriction of) a smooth function of $t$. Theorem 13 then implies that there are no singular $t$’s and that our calculations apply to all $t$. That is, while we cannot exclude singular entropy maximizers a priori, we are able to exclude them a posteriori.
We note that the analysis of optimizing the function $F$ in equation (20) is the starting point of our discussion of ERGMs in Section 6.
3. Proof of Theorem 1
We now prove a slightly more quantitative version of Theorem 1:
Theorem 14.
For each fixed $e<1/2$ and all $t$ sufficiently small, the optimal graphon with edge/triangle densities $(e,t)$ is unique and is symmetric bipodal, with parameters that vary analytically with $(e,t)$. As $t\to 0$, the increase $\Delta\mathbb{B}=\mathbb{B}(e,t)-\mathbb{B}(e,0)$ in the Boltzmann entropy scales as $t\ln(1/t)$ and the Lagrange multiplier $\beta$ scales as $\ln(1/t)$.
3.1. Strategy
The proofs of Theorems 14, 17 and 20 all follow the same general outline. We will present the proof of Theorem 14 in full detail. The subsequent proofs of Theorems 17 and 20 will be somewhat abbreviated, concentrating on what is different in those cases.
Using the fact that the unique entropy maximizing graphon $g_{0}$ at $(e,0)$ is symmetric bipodal, we show that the optimizing graphons at points near the boundary have the same general structure away from an exceptional set of small area. Specifically, we partition the unit interval into subsets $I_{1}$, $I_{2}$ and $I_{3}$ such that the columns $g_{x}$ of the optimal graphon are $L^{2}$-close to the columns of the first pode of $g_{0}$ when $x\in I_{1}$ and are $L^{2}$-close to the columns of the second pode of $g_{0}$ when $x\in I_{2}$, and where $I_{3}$ has small measure. At this stage, we do not have any control over $g_{x}$ when $x\in I_{3}$.
Knowing the columns $g_{x}$ when $x\in I_{1}\cup I_{2}$ (to within a small error in $L^{2}$) gives us pointwise control of the function $G(x,y)$ on $(I_{1}\cup I_{2})\times(I_{1}\cup I_{2})$. The Euler-Lagrange equations (16) then give us pointwise estimates of $g(x,y)$ in each of the four main rectangles.
We then study the worth functional $W(C)$. The dependence of this functional on the graphon $g$ comes via the integral $\iint C(y)C(z)g(y,z)\,dy\,dz$. Since $C$ is bounded, and since we know $g(y,z)$ away from a set of small measure, we have good control over $W(C)$. We determine that a worth-maximizing column can only take one of two approximate forms, namely those exhibited by $g_{x}$ for $x\in I_{1}$ and for $x\in I_{2}$. We then reassign each point $x\in I_{3}$ to $I_{1}$ or $I_{2}$, depending on which worth-maximizing form $g_{x}$ takes. The result is then a graphon with two (approximate) podes.
Using the pointwise equations (16), we bound the variation of $g(x,y)$ in each rectangle $I_{i}\times I_{j}$ by a multiple of the variation in a neighboring rectangle. Combining these results, the variation in each rectangle is bounded by a small multiple of itself, and so must be zero. That is, our optimal graphon must be exactly bipodal.
The space of bipodal graphons with given values of $(e,t)$ is only 2-dimensional. Using ordinary 2-dimensional calculus, we determine that the entropy $S(g)$ is maximized when the graphon is symmetric.
To account for the fact that the Lagrange multipliers $\alpha$ and $\beta$ are only defined for almost every $t$ and not necessarily for every $t$, we analyze optimal graphons in two passes. In the first pass we use the worth function and pointwise equations (16), as outlined above, to establish that the optimal graphon is symmetric bipodal for almost every $t$ that is sufficiently small. This is the bulk of the proof of Theorem 14.
This shows that the Boltzmann entropy is almost everywhere equal to the Shannon entropy of a symmetric bipodal graphon, which is a differentiable function of $t$ for each $e$. By Theorem 13, this then implies that there are no singular values of $t$ and that our arguments apply at all sufficiently small values of $t$.
The same two-pass argument applies to the proofs of Theorems 17 and 20, only replacing “symmetric bipodal” with the particular graphon symmetry described in those theorems. Specifically, in the first pass we show that the Boltzmann entropy is almost everywhere equal to the maximum Shannon entropy among multipodal graphons of a certain sort. The solution to the resulting finite-dimensional maximization problem yields a smooth function of $t$. Theorem 13 then says that our results apply at every $t$.
Remark 15.
As is standard when working with $L^{2}$ spaces, the proofs of Theorems 14, 17 and 20 are written as if our graphons were actual functions on $0,1]^{2}$. But in fact they are equivalence classes of functions that agree away from a set of measure zero. This has several consequences, none of which materially affect the flow of the proofs:
  • When we use the variational equations (16), the results apply almost everywhere, not literally everywhere. Whenever we use those equations to compute an upper or lower bound on a graphon, that bound should always be understood to mean “apart from on a set of measure zero”.
  • When we speak of the “maximum” value of a graphon $g$ on a region, we actually mean the essential supremum of the function, namely the smallest number $M$ such that $g(x,y)\leq M$ on a set of full measure. The “minimum” is similar.
  • Since we are working with functions mod sets of measure zero, we are free to change the value of our graphon on sets of measure zero whenever we wish. In this way, we could make the variational equations apply everywhere, or we could make the maximum of a function equal the essential supremum. We could, but we won’t actually subject the reader to such painstaking bookkeeping!
Instead, we will not keep track of sets of measure zero in these proofs, such as deciding whether a pode contains its endpoints. All the important properties of graphons (or at least all the properties considered in this paper) are based on integrals, for which sets of measure zero don’t matter at all.
Having explained the process in all three settings, we return our focus to the first pass.
3.2. Defining approximate podes
There is a unique entropy maximizer $g_{0}$ at $(e,0)$ on the bottom boundary of the Razborov triangle (up to measure-preserving transformations of the unit interval, as usual). This graphon is symmetric bipodal, taking values 0 on the diagonal blocks and $2e$ on the off-diagonal blocks. As we approach the bottom boundary of the Razborov triangle, we claim that any sequence $\{g_{i}\}$ of entropy maximizers must converge (after appropriate measure-preserving transformations) to $g_{0}$ in $L^{2}$.
To see this, we invoke the compactness of the space of reduced graphons in the cut metric. A subsequence must converge to a limit $g_{\infty}]$ in the cut metric. Lemma 2.1 in [14] proves that $S$ is upper-semicontinuous on $\widetilde{\mathcal{W}}$ in the cut metric $\delta_{cut}$. This implies that the limit of the entropies of the entropy maximizers is at least $S([g_{0}])$, so we must have $S([g_{\infty}])\geq S([g_{0}])$. But $g_{0}]$ is the unique entropy maximizing reduced graphon at $(e,0)$, so $g_{\infty}]=[g_{0}]$.
The entire sequence $\{[g_{i}]\}$, and not just a subsequence, must converge to $g_{0}]$. If it did not, we could find a subsequence where all points were bounded away from $g_{0}]$ in the cut metric $\delta_{cut}$. Applying the previous argument to this subsequence would then yield a contradiction. That is, after applying appropriate measure-preserving transformations of $[0,1]$, the sequence $\{g_{i}\}$ of entropy maximizing graphons must converge to $g_{0}$ in the cut distance $d_{cut}$.
Let $I_{1}=[0,1/2]$ and $I_{2}=[1/2,1]$ be the two podes of $g_{0}$. By the definition of the cut distance, the average value of $g_{i}$ must converge to 0 on $I_{1}\times I_{1}$ and $I_{2}\times I_{2}$ and to $2e$ on $I_{1}\times I_{2}$ and $I_{2}\times I_{1}$. The variance of $g_{i}$ must go to zero on each of these rectangles, or else $\lim S(g_{i})$ would be strictly less than $S(g_{0})$. Having $\lim S(g_{i})<S(g_{0})$ is impossible because there exist explicit symmetric bipodal graphons with $t\to 0$ whose entropies give a lower bound for $S(g_{i})$ and whose entropies approach $S(g_{0})$ as $t\to 0$.
Since the mean of $g_{i}$ in each rectangle approaches the value of $g_{0}$ and since the variance goes to zero, $\{g_{i}\}$ approaches $g_{0}$ in $L^{2}$. That is, for every $\epsilon>0$ there is a $\delta>0$ such that, for all $t<\delta$ and all optimal graphons $g$ at $(e,t)$, applying a measure-preserving transformation of $[0,1]$ to $g$ we have $\|g-g_{0}\|_{2}<\epsilon$. (Note that we have not assumed that the optimal graphon $g$ is unique. That will be proven in due course.)
Let $g$ be such an optimal graphon for a particular value of $(e,t)$. Then
(30) $\epsilon^{2}\geq\int_{0}^{1}dx\int_{0}^{1}dy\big{(}g(x,y)-g_{0}(x,y)\big{)}^{2},$
so
(31) $\int_{0}^{1}\big{(}g(x,y)-g_{0}(x,y)\big{)}^{2}dy<\epsilon,$
except on a set of $x$’s of measure $\epsilon$ or less. Call that exceptional set $I_{3}$. Let $I_{1}$ and $I_{2}$ be the intersection of $I_{3}^{c}$ with $[0,1/2]$ and $[1/2,1]$, respectively. Let $C_{1}$ and $C_{2}$ to be the columns of $g_{0}$ on the two podes, namely $2e$ times the indicator functions of $[1/2,1]$ and $[0,1/2]$, respectively. We have broken the unit interval into three pieces $I_{1}$, $I_{2}$, $I_{3}$, such that:
  • For all $x\in I_{1}$, $\|g_{x}-C_{1}\|_{2}<\sqrt{\epsilon}$.
  • For all $x\in I_{2}$, $\|g_{x}-C_{2}\|_{2}<\sqrt{\epsilon}$.
  • When $x\in I_{3}$ we do not yet have any estimates on $g_{x}$.
We will refer to the sets $I_{1}$, $I_{2}$ and $I_{3}$ as podes, even though we are not assuming that the graphon $g$ is exactly tripodal.
3.3. Variational equations
We now use the pointwise variation equations (16) to get some preliminary estimates on $g(x,y)$. The first two derivatives of the function $H$ are
(32) $H^{\prime}(u)=\ln(1-u)-\ln(u),\qquad H^{\prime\prime}(u)=-\left(\frac{1}{u}+\frac{1}{1-u}\right).$
The quantity $G(x,y)$ is the $L^{2}$-inner product of $g_{x}$ and $g_{y}$, which we denote $\langle g_{x}|g_{y}\rangle$. That is,
(33) $G(x,y)=\langle g_{x}|g_{y}\rangle=\int_{0}^{1}g(x,z)g(y,z)dz.$
If $x$ and $y$ are both in $I_{1}$, or both in $I_{2}$, then $G(x,y)=2e^{2}+O(\sqrt{\epsilon})$. If one is in $I_{1}$ and the other is in $I_{2}$, then $G(x,y)=O(\sqrt{\epsilon})$. If either or both are in $I_{3}$, then our estimates do not apply.
For $(x,y)\in I_{1}\times I_{1}$ or $I_{2}\times I_{2}$, we have
(34) $H^{\prime}(g(x,y))=\alpha+2\beta e^{2}(1+O(\sqrt{\epsilon}))=2\beta e^{2}(1+O(\sqrt{\epsilon})),$
so
(35) $g(x,y)=\exp(-2e^{2}\beta(1+O(\sqrt{\epsilon})).$
(Since $\beta$ is divergent as $t\to 0$ but $\alpha$ is not, we can absorb $\alpha$ into the $O(\beta\sqrt{\epsilon})$ error.) This means that the contribution of $g(x,y)$ in $I_{1}\times I_{1}$ or $I_{2}\times I_{2}$ to $\beta G$ goes as $\beta$ times a negative exponential in $\beta$, and thus has an extremely small effect on the value of $g(x,y)$ in $I_{1}\times I_{2}$ or $I_{2}\times I_{1}$.
However, we cannot yet precisely estimate $g(x,y)$ in those regions because $G(x,y)=\langle g_{x}|g_{y}\rangle$ also gets a contribution, potentially of order $\epsilon$, from $z\in I_{3}$.
3.4. Maximizing worth and eliminating $I_{3}$
Let $C:[0,1]\to[0,1]$ be a function whose worth we wish to estimate. Let
(36) $a=2\int_{0}^{1/2}C(y)\,dy,\qquad b=2\int_{1/2}^{1}C(y)\,dy.$
That is, $a$ and $b$ are the average values of $C$ on $[0,1/2]$ and $[1/2,1]$.
We now consider the three expressions that contribute to $W(C)$:
  • The entropy term $2\int_{0}^{1}H(C(y))\,dy$ is bounded above by $H(a)+H(b)$, thanks to $H^{\prime\prime}$ being everywhere negative.
  • The term $2\alpha\int_{0}^{1}C(y)\,dy$ is exactly $\alpha(a+b)$.
  • The term $\beta\iint C(y)C(z)g(y,z)\,dy\,dz$ is approximately $e\beta ab$.
Recall that the worth of a column is
(37) $W(C)=2\int_{0}^{1}H(C(y))\,dy-2\alpha\int_{0}^{1}C(y)\,dy-\beta\iint C(y)C(z)g(y,z)\,dy\,dz.$
If $g$ were equal to $g_{0}$, maximizing $W(C)$ would involve taking $C(y)$ to be constant on $[0,1/2]$ and constant on $[1/2,1]$ and choosing $a$ and $b$ to maximize
(38) $H(a)+H(b)-\alpha(a+b)-e\beta ab.$
(Because of the small differences between $g$ and $g_{0}$, this procedure only gives approximate worth maximizers, but that is enough for our purposes.)
Setting the derivatives of (38) to zero gives the equations
(39) $H^{\prime}(a)=\alpha+\beta eb,\qquad H^{\prime}(b)=\alpha+\beta ea.$
Since there is a worth-maximizer with $a$ close to 0 and $b$ close to $2e$ (namely any column with $x\in I_{1}$), and another worth-maximizer with $a$ close to $2e$ and $b$ close to 0, $\alpha$ must be close to $H^{\prime}(2e)$, while $\beta$ is large and positive.
If $a$ is substantially nonzero (say, bigger than $1/\sqrt{\beta}$), then $e\beta a$ is gigantic and $b$ is extremely close to zero, being $O(\exp(-\sqrt{\beta}))$. This makes $e\beta b$ tiny so $H^{\prime}(a)\approx\alpha\approx H^{\prime}(2e)$ and $a\approx 2e$. Similarly, if $b$ is substantially nonzero then $a$ is tiny and $b\approx 2e$. In both those cases, $W(C)\approx H(2e)-2eH^{\prime}(2e)=-\ln(1-2e)$. The third possibility is that $a$ and $b$ are both tiny, but in that case $W(C)\approx 0$, which is strictly less than $\ln(1-2e)$.
The upshot is that there are three stationary points of (38) but only two maxima, one that resembles $g_{x}$ for $x\in I_{1}$ and one that resembles $g_{x}$ for $x\in I_{2}$. Since every column $g_{x}$ with $x\in I_{3}$ must be a worth-maximizer, and since every worth-maximizer must come close to maximizing (38), every column $g_{x}$ with $x\in I_{3}$ is close in $L^{2}$ to the columns for $x\in I_{1}$ or $I_{2}$. We can then reassign the points of $I_{3}$ to $I_{1}$ or $I_{2}$ depending on the nature of $g_{x}$.
3.5. Exact bipodality
Our next step is to upgrade our $L^{2}$ estimates on the forms of the different columns into pointwise estimates. Thanks to each column of $g$ being $L^{2}$-close to a column of $g_{0}$, the function $G(x,y)=\langle g_{x}|g_{y}\rangle$ is pointwise close to $2e^{2}$ on $I_{1}\times I_{1}$ and on $I_{2}\times I_{2}$. By (16), this forces $g(x,y)$ to be exponentially small (specifically, $\exp(-\Theta(\beta)))$ in these quadrants. This in turn makes $G(x,y)$ exponentially small on $I_{1}\times I_{2}$ and $I_{2}\times I_{1}$, which means that $H^{\prime}(g)$ is exponentially close to $\alpha$ in these quadrants, and therefore that $g(x,y)$ is pointwise close to constant in these rectangles.
We next show that the optimal graphon $g$ is exactly constant on each of those rectangles. Let $A$, $B$, and $D$ be the average values of $g(x,y)$ on $I_{1}\times I_{1}$, $I_{2}\times I_{2}$ and $I_{1}\times I_{2}$, respectively. Let $\Delta_{A}$, $\Delta_{B}$ and $\Delta_{D}$ be the difference between the maximum and minimum values of $g(x,y)$ on those rectangles. Let $c$ be the width of $I_{1}$.
On $I_{1}\times I_{1}$, the quantity $G(x,y)$ is bounded below by
(40) $c(A-\Delta_{A})^{2}+(1-c)(D-\Delta_{D})^{2}$
and bounded above by
(41) $c(A+\Delta_{A})^{2}+(1-c)(D+\Delta_{D})^{2}.$
The difference between these two expressions is $4cA\Delta_{A}+4(1-c)D\Delta_{D}$.
All points satisfy the variational equations
(42) $H^{\prime}(g(x,y))=\alpha+\beta G(x,y).$
Subtracting this equation at the smallest value of $G(x,y)$ from that at the largest value, applying the mean value theorem to the left hand side, and applying our bounds on the variation in $G(x,y)$, we obtain
(43) $H^{\prime\prime}(A_{0})\Delta_{A}\leq 4Ac\beta\Delta_{A}+4D(1-c)\beta\Delta_{D},$
where $A_{0}$ is some number between $A+\Delta_{A}$ and $A-\Delta_{A}$. A little algebra then shows that
(44) $\Delta_{A}\leq\frac{4D(1-c)\beta}{-H^{\prime\prime}(A_{0})-4Ac\beta}\Delta_{D}\leq\frac{3D\beta}{-H^{\prime\prime}(A)}\Delta_{D},$
where we have used the difference between 3 and $4(1-c)\approx 2$ to cover for simplifying the denominator and replacing $A_{0}$ with $A$. A similar analysis on $I_{2}\times I_{2}$ shows that
(45) $\Delta_{B}\leq\frac{3D\beta}{-H^{\prime\prime}(B)}\Delta_{D}.$
Meanwhile, on $I_{1}\times I_{2}$, $G(x,y)$ is bounded above and below by
(46) $c(A\pm\Delta_{A})(D\pm\Delta_{D})+(1-c)(B\pm\Delta_{B})(D\pm\Delta_{D}),$
where the plus signs give an upper bound and the minus signs give a lower bound. The difference between the upper and lower bounds is
(47) $2\beta[(cA+(1-c)B)\Delta_{D}+(c\Delta_{A}+(1-c)\Delta_{B})D].$
This implies that
(48) $H^{\prime\prime}(D_{0})\Delta D\leq 2\beta(cA+(1-c)B)\Delta_{D}+2\beta D(c\Delta_{A}+(1-c)\Delta_{B}).$
A little algebra then gives
(49) $\Delta_{D}$ $\leq$ $\frac{2\beta D(C\Delta A+(1-c)\Delta B)}{-H^{\prime\prime}(D_{0})-2\beta(cA+(1-c)B}$
(50) $\leq$ $\frac{3\beta D(\Delta_{A}+\Delta_{B})}{-2H^{\prime\prime}(D)}$
(51) $\leq$ $\frac{9\beta^{2}D^{2}}{-2H^{\prime\prime}(D)}\left(\frac{-1}{H^{\prime\prime}(A)}+\frac{-1}{H^{\prime\prime}(B)}\right)\Delta_{D}.$
Now recall that $A$ and $B$ are exponentially small in $\beta$ and that
(52) $\frac{-1}{H^{\prime\prime}(A)}=A(1-A)<A\qquad\hbox{and}\qquad\frac{-1}{H^{\prime\prime}(B)}=B(1-B)<B.$
The coefficient of $\Delta_{D}$ on the right hand side of the last line goes to zero roughly as $\beta^{2}\exp(-2e^{2}\beta)$ as $t\to 0$ and $\beta\to\infty$. Once the coefficient is less than one, the only solution is $\Delta_{D}=0$, which then implies that $\Delta_{A}=0$ and $\Delta_{B}=0$. In other words, our optimal graphon is exactly bipodal.
3.6. Symmetric bipodality
All that remains is showing that the best bipodal graphon is symmetric, with pode sizes $\frac{1}{2}$ and $\frac{1}{2}$ and with $A=B$. This requires extensive calculations but no sophisticated analysis. Ultimately, it is just a (grungy) problem in multivariable calculus as follows.
For each triple $(e,t,c)$ we consider the bipodal graphon that maximizes the entropy, subject to the constraints that the edge and triangle densities are $(e,t)$ and that the first pode has width $c$. Let $S(e,t,c)$ be the entropy of this optimal graphon. We must show that this entropy is maximized at $c=1/2$. Note that this function is analytic in $c$ for fixed $(e,t)$, insofar as the parameters are determined by analytic Euler-Lagrange equations, and is even in $\Delta c:=c-\frac{1}{2}$.
When $t=0$, the function is easy to compute. The graphon must be zero on $I_{1}\times I_{1}$ and $I_{2}\times I_{2}$ and take on the constant value $\frac{e}{2c(1-c)}=\frac{2e}{1-4\Delta c^{2}}$ on $I_{1}\times I_{2}$. The entropy is then
(53) $S(e,0,c)$ $=$ $\frac{1}{2}(1-4\Delta c^{2})H\left(\frac{2e}{1-4\Delta c^{2}}\right)$
(54) $=$ $S(e,0,1/2)+2\ln(1-2e)\Delta c^{2}+O(\Delta c^{4}).$
That is, there is an entropy cost proportional to $\Delta c^{2}$ associated with having $\Delta c\neq 0$.
Now consider the effect of having $t$ nonzero. Having the graphon nonzero on $I_{1}\times I_{1}$ and $I_{2}\times I_{2}$ provides additional entropy of order $t\ln(1/t)$. Shifting the value of the graphon on $I_{1}\times I_{2}$ by an $O(t)$ amount changes the entropy by an additional $O(t)$, but since this is small compared to $t\ln(1/t)$, S(e,t,c)S(e,0,c)S(e,t,c)-S(e,0,c) is still $O(t\ln(1/t))$. In order to overcome the $2\ln(1-2e)\Delta c^{2}$ cost, we must have $\Delta c=O(\sqrt{t\ln(1/t)})$. Since $t\sim\exp(-2e^{2}\beta)$, this means that $\Delta c$ must be exponentially small in $\beta$ and in particular that $\beta\Delta c$ is a small parameter.
We now compute the quantity $G(x,y)$ in each rectangle and look at the Euler-Lagrange equations for a particular value of $\beta$:
(55) $H^{\prime}(A)$ $=$ $\alpha+\frac{\beta}{2}(A^{2}+D^{2})-\beta\Delta c(D^{2}-A^{2})$
(56) $\approx$ $\alpha+\frac{\beta}{2}D^{2}-\beta\Delta cD^{2},$
(57) $H^{\prime}(B)$ $=$ $\alpha+\frac{\beta}{2}(B^{2}+D^{2})+\beta\Delta c(D^{2}-B^{2})$
(58) $\approx$ $\alpha+\frac{\beta}{2}D^{2}+\beta\Delta cD^{2},$
(59) $H^{\prime}(D)$ $=$ $\alpha+\frac{\beta D}{2}(A+B+2\Delta c(A-B))$
(60) $\approx$ $\alpha,$
where in our approximations we use the fact that $A$ and $B$ are exponentially small in $\beta$. Since $D\approx 2e$, this makes $\alpha\approx H^{\prime}(2e)$. The terms proportional to $\Delta c$ serve to multiply $A$ by a factor of $\exp(-4e^{2}\beta\Delta c)\approx 1-4e^{2}\beta\Delta c$ and to multiply $B$ by a factor of $\exp(4e^{2}\beta\Delta c)\approx 1+4e^{2}\beta\Delta c$. These changes in the values of $A$ and $B$ (relative to their values when $\Delta c=0$) slightly change the triangle density for a given value of $\beta$, but only by a fraction $O(\beta\Delta c^{2})$. Likewise, the contribution to the entropy of the $I_{1}\times I_{1}$ and $I_{2}\times I_{2}$ squares changes by a fraction $O(\beta\Delta c^{2})$. However, that entropy is only $O(t\ln(1/t))$, so we are dealing with an expression that is
(61) $O(\beta t\ln(1/t)\Delta c^{2})=O(t(\ln(1/t)^{2})\Delta c^{2}),$
since $\beta=O(\ln(1/t))$. This possible entropy gain from having $\Delta c\neq 0$ is much smaller than the $2\ln(1-2e)\Delta c^{2}$ cost, so the optimal value of $\Delta c$ is exactly zero. That is, we must have $c=1/2$.
When $c=1/2$, two of the Euler-Lagrange equations read:
(62) $H^{\prime}(A)$ $=$ $\alpha+\beta(A^{2}+D^{2})/2,$
(63) $H^{\prime}(B)$ $=$ $\alpha+\beta(B^{2}+D^{2})/2.$
If $A>B$, then the right hand side of the first equation is greater than that of the second, so $H^{\prime}(A)>H^{\prime}(B)$. But that is a contradiction, since $H^{\prime}(u)=\ln(1-u)-\ln(u)$ is a decreasing function of $u$. Likewise, we cannot have $A<B$. So $A$ and $B$ must be equal, making our optimal graphon symmetric bipodal.
The parameters of a symmetric bipodal graphon are uniquely (and analytically) determined by $(e,t)$.
We also consider how various quantities scale as $t\to 0$. After setting $c=1/2$ and $B=A$, a direct calculation shows that
(64) $t=\frac{3}{4}AD^{2}+\frac{1}{4}A^{3},$
so
(65) $A=\frac{4t}{3D^{2}}+O(t^{3})=\frac{t}{3e^{2}}+O(t^{2}),$
where we have used the fact that D=2eAD=2e-A. Since $A$ was exponentially small in $\beta$, $\beta$ must scale as $\ln(1/t)$. The entropy is
(66) $\frac{1}{2}(H(A)+H(2e-A))=\frac{1}{2}H(2e)-\frac{1}{2}A\ln(A)+O(A),$
so $S(g)-\frac{1}{2}H(2e)$ scales as $t\ln(1/t)$.
The Boltzmann entropy $\mathbb{B}(e,t)$ is equal to the Shannon entropy $S(g)$ of the optimal graphon at $(e,t)$, so $\Delta\mathbb{B}:=\mathbb{B}(e,t)-\mathbb{B}(e,0)=S(g)-\frac{1}{2}H(2e)$. Since $A\approx(t/3e^{2})$ and $A\approx\exp(-2e^{2}\beta)$, $\beta\approx\frac{-1}{2e^{2}}\ln(t/3e^{2})\sim\ln(1/t)$. This completes the proof of Theorem 14. ∎
Note that we have actually proved something slightly stronger than Theorem 14. We started with an optimal graphon, applied measure preserving transformations of $[0,1]$, and wound up with a symmetric bipodal graphon. That is, the optimal graphon in $\mathcal{W}_{e,t}$ is unique up to group equivalence. However, a reduced graphon in $\widetilde{\mathcal{W}}_{e,t}$ is a weak equivalence class, with all representatives of this class being constrained entropy maximizers. We thus conclude that
Corollary 16.
For all $e<1/2$ and for all $t$ small enough that Theorem 14 applies, any graphon $g\in\mathcal{W}_{e,t}$ that is weakly equivalent to a symmetric bipodal graphon is group equivalent to a symmetric bipodal graphon.
4. Proof of Theorem 2
As in the last section, we will prove a slightly extended version of Theorem 2:
Theorem 17.
Let $n\geq 1$ be an integer. For every $e\in\left(\frac{n}{n+1},\frac{n+1}{n+2}\right)$, with corresponding minimal triangle density $t_{0}$ (depending on $e$), and for all $\Delta t$ sufficiently small, the optimal graphon with edge/triangle densities $(e,t_{0}+\Delta t)$ is unique and $n+2$-podal, with $(n,2)$ symmetry. Asymptotically, $\Delta\mathbb{B}=\mathbb{B}(e,t)-\mathbb{B}(e,t_{0})$ scales as $\sqrt{\Delta t}$. and the Lagrange multiplier $\beta$ scales as $1/\sqrt{\Delta t}$. In the optimal graphon, the diagonal entries are all $\exp(-\Theta(\beta))$ and, except for the $(n+1,n+2)$ and $(n+2,n+1)$ entries, the off-diagonal entries are all $1-\exp(-\Theta(\beta))$.
Proof.
The proof of Theorem 17 (and therefore Theorem 2) follows the same script as the proof of Theorem 14, namely
  1. (1)
    Using the proximity of an entropy-maximizing graphon $g$ at $(e,t)$ to the unique entropy-maximizing graphon $g_{0}$ at $(e,t_{0})$ to define approximate podes $I_{1},\ldots,I_{n+3}$ where the columns with $x\in I_{j}$ with $j\leq n+2$ are $L^{2}$-close to the corresponding columns of $g_{0}$, and where the exceptional set $I_{n+3}$ is small.
  2. (2)
    Using the Euler-Lagrange equations to show that all of the graphon values are exponentially close to 0 or 1, except on $I_{n+1}\times I_{n+2}$, $I_{n+2}\times I_{n+1}$, or when one of the coordinates is in $I_{n+3}$.
  3. (3)
    Showing that the only possible worth-maximizing columns are small perturbations of the columns of $g_{0}$, thus allowing us to reassign the points of $I_{n+3}$ to the other podes.
  4. (4)
    Bounding the variation in $g(x,y)$ in each rectangle $I_{i}\times I_{j}$ by a small multiple of the variation in other rectangles. Combining estimates, this shows that the variation in each rectangle is bounded by a small multiple of itself, and must therefore be zero.
  5. (5)
    Analyzing the finite-dimensional space of $(n+2)$-podal graphons near $g_{0}$ and determining that the best one has $(n,2)$ symmetry. We then determine how $S$, $\beta$, and various entries of the optimal graphon scale with $\Delta t$.
  6. (6)
    On the first pass, steps (1–5) only apply at values of $t$ for which the Lagrange multipliers $\alpha$ and $\beta$ are well-defined and finite. Extending the results to all sufficiently small values of $\Delta t$ then follows from Theorem 13.
There is one important difference between the situation of Theorem 14 and that of Theorem 17. The additional podes that appear on the scallops provide an additional, and more efficient, means of generating entropy at the expense of added triangles. As a result, $\Delta\mathbb{B}$ scales as $\sqrt{\Delta t}$ rather than $\Delta t\ln(1/\Delta t)$. Before getting into the details of the proof, we explain how this works, starting near the first scallop, with $e\in(\frac{1}{2},\frac{2}{3})$.
Consider tripodal graphons of the form shown in Figure 4. The total edge density is
(67) $e=2c(1-c)+\frac{1}{2}p(1-c)^{2},$
so we must have
(68) $p=\frac{2(e-2c(1-c))}{(1-c)^{2}}.$
The triangle density is
(69) $t=\frac{3}{2}pc(1-c)^{2}=3c(e-2c(1-c))=3ec-6c^{2}+6c^{3}.$
Taking derivatives, we see that
(70) $\frac{dt}{dc}=3(6c^{2}-4c+e)\hbox{ and }\frac{d^{2}t}{dc^{2}}=36c-12.$
The first derivative is zero when
(71) $c=\frac{1}{3}\left(1+\sqrt{1-\frac{3e}{2}}\right).$
Since $d^{2}t/dc^{2}$ is always positive, this gives the minimum triangle density among graphons of this kind. In fact, it minimizes $t$ among all possible graphons [40] and is the unique optimal graphon with densities $(e,t_{0})$ [42].
Figure 4. A tripodal graphon of the form seen on the first scallop
Now imagine varying $c$ and $p$ while preserving the structure of Figure 4. The entropy of the graphon displayed in Figure 4 is
(72) $S=\frac{1}{2}(1-c)^{2}H(p),$
where $p$ is given by equation (68). A little algebra then gives
(73) $p=4-4(1-c)^{-1}+2e(1-c)^{-2},$
so
(74) $\frac{dp}{dc}=4e(1-c)^{-3}-4(1-c)^{-2}=\frac{4(e-(1-c))}{(1-c)^{3}}.$
We then compute
(75) $\frac{dS}{dc}$ $=$ $-(1-c)H(p)+\frac{1}{2}(1-c)^{2}H^{\prime}(p)\frac{dp}{dc}$
(76) $=$ $-(1-c)H(p)+\frac{2H^{\prime}(p)(e+c-1)}{1-c}$
(77) $=$ $-(1-c)H(p)+\frac{H^{\prime}(p)}{1-c}(p(1-c)^{2}+6c-4c^{2}-2)$
(78) $=$ $(1-c)(pH^{\prime}(p)-H(p))+(4c-2)H^{\prime}(p)$
(79) $=$ $(1-c)\ln(1-p)+4c-2)(\ln(1-p)-\ln(p))$
(80) $=$ $(3c-1)\ln(1-p)+2(1-2c)\ln(p).$
Since $c$ is between $\frac{1}{2}$ and $\frac{2}{3}$, the coefficients of $\ln(1-p)$ and $\ln(p)$ are both positive, making $\frac{dS}{dc}$ negative. We can increase the entropy to first order by decreasing $c$. That only increases the triangle count to second order in $\Delta c$, so we have achieved an entropy increase that scales as the square root of $\Delta t$.
The situation is similar near the other scallops. There is a family of graphons parametrized by the size $c$ of each of the $n$ identical podes, as in Figure 5 for $n=3$. There is a value $c_{0}$ that minimizes the triangle density, but $dS/dc$ is not zero at $c=c_{0}$. Instead, the calculation shown in the next paragraph shows that $dS/dc$ is negative for all relevant values of $c$. As a result, we can increase $S$ to first order in $\Delta c$ by decreasing $c$ while only increasing $t$ to second order, so $\Delta S\sim\sqrt{\Delta t}$.
Figure 5. A multipodal graphon of the form seen on the scallops, in this case with $n=3$
For a general value of $n$, the graphon on the $n$-th scallop is 1 everywhere except on the diagonal blocks and on the two off-diagonal blocks in the upper right corner. The edge density is
(81) $e=1-nc^{2}-(1-nc)^{2}+\frac{(1-nc)^{2}}{2}p.$
This means that
(82) $p$ $=$ $\frac{2}{(1-nc)^{2}}\left(e+nc^{2}-1+(1-nc)^{2}\right)$
(83) $=$ $2\left(e-\frac{n-1}{n}\right)(1-nc)^{-2}-\frac{4}{n}(1-nc)^{-1}+\frac{2(n+1)}{n}.$
Taking a derivative with respect to $c$ is then easy:
(84) $\frac{dp}{dc}=\frac{4n}{(1-nc)^{3}}(e+c-1).$
The entropy is $S=\frac{(1-nc)^{2}}{2}H(p)$ and derivative of $S$ with respect to $c$ is
(85) $\frac{dS}{dc}$ $=$ $-n(1-nc)H(p)+\frac{1}{2}(1-nc)^{2}H^{\prime}(p)\frac{dp}{dc}$
(86) $=$ $-n(1-nc)H(p)+\frac{2nH^{\prime}(p)(e+c-1)}{1-nc}$
(87) $=$ $-n(1-nc)H(p)+nH^{\prime}(p)((n+1)c-1+p)$
(88) $=$ $n(1-nc)(pH^{\prime}(p)-H(p))+2n((n+1)c-1)H^{\prime}(p)$
(89) $=$ $-n(1-nc)\ln(1-p)+2n((n+1)c-1)(\ln(1-p)-\ln(p))$
(90) $=$ $n((n+2)c-1)\ln(1-p)+2n(1-(n+1)c)\ln(p).$
Since $c$ is between $\frac{1}{n+2}$ and $\frac{1}{n+1}$, the coefficients of $\ln(p)$ and $\ln(1-p)$ are both positive, making each term negative, so $\frac{ds}{dc}<0$, as claimed.
4.1. Defining approximate podes
We now turn to the details of the proof. As usual, let $g_{0}$ be the unique entropy maximizer $g_{0}$ at $(e,t_{0})$. This graphon takes the form shown in Figure 5, with
(91) $c=c_{0}=\frac{1+\sqrt{1-\frac{n+2}{n+1}e}}{n+2},$
which is the value of $c$ that minimizes
(92) $t=n(n+1)(n+2)c^{3}-3n(n+1)c^{2}+3nec.$
As we approach the scallop, any sequence of entropy maximizers must converge (after applying measure preserving transformations of $[0,1]$) to $g_{0}$ in $L^{2}$ by exactly the same argument as in the proof of Theorem 14.
As before, any sequence of entropy maximizing reduced graphons must converge to $g_{0}]$ in the cut metric, implying than any sequence of entropy maximizing graphons must are equivalent to graphons $g_{i}$ that converge to $g_{0}$ in the cut distance. By the definition of the cut distance, the average value of $g_{i}$ on each rectangle defined by the podes of $g_{0}$ must approach the (constant) value of $g_{0}$ on that rectangle. Since $\lim S(g_{i})=S(g_{0})$, the variance of $g_{i}$ must go to zero on each of these rectangles, so the graphons $g_{i}$ converge to $g_{0}$ in $L^{2}$.
We pick a sufficiently small value of $\epsilon$ and consider values of $t$ small enough that $\|g-g_{0}\|_{L^{2}}<\epsilon$ for each optimal graphon $g$. Let $I_{1}$, $I_{2}$, $\ldots$, $I_{n+2}$ be the subsets of the podes of $g_{0}$ for which $g_{x}$ lies within $\sqrt{\epsilon}$ in $L^{2}$ of the corresponding column of $g_{0}$ and let $I_{n+3}$ be the exceptional set where $g_{x}$ is not close to the corresponding column of $g_{0}$. Note that $I_{n+3}$ may contain points $x$ where $g_{x}$ is close to a different column of $g_{0}$. Those points will soon be reassigned.
4.2. Variational equations
Next we need to compute $G(x,y)$ in different cases. Let $G_{i,j}$ denote a typical value of $G(x,y)$ when $x\in I_{i}$ and $y\in I_{j}$. We can estimate these quantities to within $O(\sqrt{\epsilon})$ using the columns of $g_{0}$. Thanks to our $(n,2)$ symmetry, there are only five different numbers to compute, namely $G_{1,1}$, $G_{1,2}$, $G_{1,{n+1}}$, $G_{n+1,n+1}$ and $G_{n+1,n+2}$. The results are
(93) $G_{1,1}$ $\approx$ $1-c,$
(94) $G_{1,2}$ $\approx$ $1-2c,$
(95) $G_{1,n+1}$ $\approx$ $(n-1)c+\frac{1-nc}{2}p,$
(96) $G_{n+1,n+1}$ $\approx$ $nc+\frac{1-nc}{2}p^{2},$
(97) $G_{n+1,n+2}$ $\approx$ $nc,$
where $\approx$” means “equal to within $O(\sqrt{\epsilon})$”. Note that $G_{1,1}$ and $G_{n+1,n+1}$ are greater than $G_{n+1,n+2}$ by amounts that are $\Omega(1)$ as $t\to 0$ while $G_{1,2}$, and $G_{1,n+1}$ are less than $G_{n+1,n+2}$ by amounts that are $\Omega(1)$. Multiplying by $\beta$ and adding $\alpha$, and using the fact that
(98) $H^{\prime}(p)=\alpha+\beta G_{n+1,n+2}=(\alpha+nc)+O(\beta\sqrt{\alpha}).$
we get that $H^{\prime}(g(x,y))$ is $\Omega(\beta)$ on the diagonal rectangles that do not involve $I_{n+3}$ and is $\Omega(\beta)$ on the off-diagonal blocks that do not involve $I_{n+3}$, with the exception of $I_{n+1}\times I_{n+2}$ and $I_{n+2}\times I_{n+1}$. This implies that $g(x,y)$ is exponentially small (that is, $\exp(-\Omega(\beta))$) on the diagonal blocks, exponentially close to 1 on all but two of the off-diagonal blocks, and of course is close to $p$ on $I_{n+1}\times I_{n_{2}}$ and $I_{n+2}\times I_{n+1}$.
4.3. Maximizing worth
Let $C:[0,1]\to[0,1]$ be a function whose worth we aim to maximize. For each $i=1,2,\ldots,n+2$, let $a_{n+2}$ be the average values of $C(y)$ for $y\in I_{i}$. Of the terms contributing to $W(C)$, the entropy term is bounded above by $2c\sum_{i=1}^{n}H(a_{i})+(1-nc)(H(a_{n+1})+H(a_{n+2}))$, since fluctuations in $C$ within each pode can only decrease the entropy. The edge density term is $\alpha(2c\sum_{i=1}^{n}a_{i}+(1-nc)(a_{n+1}+a_{n+2}))$.
The most important term comes from triangles. To within the accuracy of our approximation that $g(y,z)$ is constant on each rectangle, it is the quadratic function
(99) $\beta\sum_{i,j=1}^{n+2}M_{ij}a_{i}a_{j},$
where $M_{ij}$ is the integral of $g(y,z)$ over $I_{i}\times I_{j}$.
If we are at a maximum of $W$, then the gradient of $W$ must be zero and the Hessian must be negative semi-definite. The Hessian of $W$ with respect to the variables $\{a_{i}\}$ is precisely $2\beta M$ plus diagonal terms proportional to $H^{\prime\prime}(a_{i})$. The matrix $M$ is (nearly) zero on the diagonal, with all of the off-diagonal terms being close to 1 or $p$, and so has eigenvalues of both signs. The only way for the Hessian to be negative semi-definite is for all but one of entries $H^{\prime\prime}(a_{i})$ to be at least of order $\beta$. In other words, all columns that maximize $W$ must have every entry but one (or every entry) approximately equal to 0 or 1. In terms of $G$, for $y$ in any pode but one, |G(x,y)nc||G(x,y)-nc| must exceed $\Theta(1/\beta)$.
We now examine the possibilities.
  • If $G(x,y)\approx nc$ for $y\in I_{1}$, then $2c\sum_{i=2}^{n}a_{i}+(1-nc)(a_{n+1}+a+{n+2})\approx 2nc$. But that is impossible if each $a_{i}$ (other than $a_{1}$) is equal to 0 or 1. Contradiction. Likewise, it is not possible to have $G(x,y)\approx nc$ for $y\in I_{2},\ldots,I_{n}$. The first $n$ variables $a_{i}$ are all either pegged to 0 or to 1.
  • If two or more of the variables $a_{1},\ldots,a_{n}$ are pegged to 0, then $G(x,y)<nc$ for all $y$, so $g(x,y)\approx 1$ for all $y$, which is a contradiction. Thus either one or none of the first $n$ $a_{i}$’s is pegged to 0 and the rest are pegged to 1.
  • If exactly one of these variables is pegged to 0, then for $y\in I_{n+1}$ or $y\in I_{n+2}$ we have $G(x,y)\leq p\frac{1-c}{2}+(n-1)c<nc$, so $a_{n+1}$ and $a_{n+2}$ are pegged to 1. In other words, our column is just like the columns when $x$ is in one of the first $n$ podes.
  • If all of the variables $a_{1},\ldots,a_{n}$ are pegged to 1, then we examine $a_{n+1}$ and $a_{n+2}$. Neither one is pegged to 1, since $G(x,y)$ is at least $nc$ for $y$ in either $I_{n+1}$ or $I_{n+2}$. They cannot both be pegged to 0, since that would make $G(x,y)=nc$ in both podes, meaning that the values are not pegged and the Hessian is not negative-definite. Thus one value must be pegged to 0 while the other is intermediate between 0 and 1. The closeness of all but one $a_{i}$ to 0 or 1 gives us the same equation for the remaining $a_{i}$ as satisfied by the actual columns of $I_{n+1}$ or $I_{n+1}$, implying that the final $a_{i}$ must be close to $p$. That is, $C$ is close to the actual columns when $x\in I_{n+1}$ or $I_{n+2}$.
The upshot is that all worth-maximizers are already $L^{2}$-close to columns of $g_{0}$. Since each column is a worth-maximizer, we can reassign all of the points of $I_{n+3}$ to other podes. Note that this reassignment can result in the podes of $g$ having sizes that are slightly different from those of the corresponding podes of $g_{0}$.
Controlling the columns to within a small $L^{2}$ error gives us pointwise control over $G(x,y)=\langle g_{x}|g_{y}\rangle$. On all rectangles except for $I_{n+1}\times I_{n+2}$ and $I_{n+2}\times I_{n+1}$, this forces $g(x,y)$ to be exponentially close to 0 or 1. This makes $G(x,y)$ exponentially close to constant on $I_{n+1}\times I_{n+2}$ and $I_{n+2}\times I_{n+1}$ and so makes $g(x,y)$ exponentially close to a constant (that is close to $p$, but not necessarily exponentially close) on these rectangles.
4.4. Exact multipodality
So far we have shown that an optimal graphon has to be approximately multipodal. There are $n$ podes $I_{1},\ldots,I_{n}$ of width close to
(100) $c=\frac{1+\sqrt{1-\frac{n+2}{n+1}e}}{n+2}$
and two podes of width close to $\frac{1-nc}{2}$. The graphon is exponentially close to 0 on the diagonal blocks, exponentially close to 1 on all of the off-diagonal blocks but two, and close to $p$ on $I_{n+1}\times I_{n+2}$ and $I_{n+2}\times I_{n+1}$. We next show that the graphon is exactly constant on each rectangle. The proof is essentially a rerun of the analogous step in the proof of Theorem 14, only with more terms.
Let $g_{ij}$ be the average value of the graphon on the rectangle $I_{i}\times I_{j}$ and let $\Delta g_{ij}$ be the difference between the greatest and lowest value of the graphon in that rectangle. Let $c_{i}$ be the width of $I_{i}$. We have already determined that all $g_{ij}$’s except for $g_{n+1,n+2}$ and $g_{n+2,n+1}$ are exponentially close (in $\beta$) to 0 or 1, and hence that $1/H^{\prime\prime}(g_{ij})$ is exponentially small.
If $x\in I_{i}$ and $y\in I_{j}$, then the maximum and minimum possible values of $G(x,y)$, and their difference, are
(101) max $=$ $\sum_{k=1}^{n+2}c_{k}(g_{ik}+\Delta g_{ik})(g_{jk}+\Delta g_{jk}),$
(102) min $=$ $\sum_{k=1}^{n+2}c_{k}(g_{ik}-\Delta g_{ik})(g_{jk}-\Delta g_{jk}),$
(103) difference $=$ $\sum_{k=1}^{n+2}2c_{k}(g_{ik}\Delta g_{jk}+g_{jk}\Delta g_{ik}).$
Applying the mean value theorem to the Euler-Lagrange equations, and noting that $H^{\prime\prime}(u)$ is always negative, we have
(104) $H^{\prime\prime}(g_{ij,0})\Delta g_{ij}\leq 2\beta\sum_{k=1}^{n+2}c_{k}(g_{ik}\Delta g_{jk}+g_{jk}\Delta g_{ik}),$
where $g_{ij,0}$ is some number between the values of $g$ corresponding to the maximum and minimum possible values of $G(x,y)$.
The sum on the right contains terms proportional to $\Delta g_{ij}$ itself, coming from $k=i$ or $k=j$. We bring those terms to the left hand side, noting that coefficients of those terms are much smaller than $H^{\prime\prime}(g_{ij,0})$. When $\{i,j\}\neq\{n+1,n+2\}$, this is because $g_{ij,0}$ is exponentially close to 0 or 1, so $H^{\prime\prime}(g_{ij,0})$ is exponentially large, while the coefficients on the right hand side are $O(\beta)$. When $\{i,j\}=\{n+1,n+2\}$, this is because the coefficients of $\Delta g_{n+1,n+2}$ on the right hand side are proportional to $\beta g_{n+1,n+1}$ or $\beta g_{n+2,n+2}$, both of which are exponentially small. By changing the factor of 2 on the right hand side to a 3, we can absorb these small corrections to the coefficient of $\Delta g_{ij}$ and also replace $H^{\prime\prime}(g_{ij,0})$ with just $H^{\prime\prime}(g_{ij})$. We also bound $g_{ik}$ and $g_{jk}$ by 1. The upshot is that
(105) $\Delta g_{ij}$ $\leq$ $\frac{3\beta}{-H^{\prime\prime}(g_{ij})}\sum_{k}c_{k}(\Delta g_{jk}+\Delta g_{ik})$
(106) $\leq$ $\frac{6\beta}{-H^{\prime\prime}(g_{ij})}\max(\Delta g_{ik}\hbox{ or }\Delta g_{jk}),$
where the sum on the first line and the maximum on the second line skips terms involving $\Delta g_{ij}$ itself.
Whenever $\{i,j\}\neq\{n+1,n+2\}$, $\frac{6\beta}{-H^{\prime\prime}(g_{ij})}$ is exponentially small, so $\Delta g_{ij}$ is bounded by a tiny multiple of a sum of similar errors. In particular, the largest $\Delta g_{ij}$ of this sort is bounded by a sum of contributions much smaller than itself, possibly plus a contribution from $\Delta g_{n+1,n+2}$. The conclusion is that all $\Delta g_{ij}$’s other than $\Delta g_{n+1,n+2}$ are bounded by $\beta\exp(-\Omega(\beta))\Delta g_{n+1,n+2}$.
Now consider the equation for $\Delta g_{n+1,n+2}$. This equation indicates that $\Delta g_{n+1,n+2}$ is bounded by an $O(1)$ multiple of $\beta$ times the largest of the remaining $\Delta g_{ij}$’s, and so is bounded by a constant times $\beta^{2}\exp(-\Omega(\beta))\Delta g_{n+1,n+2}$. When $\beta$ is large, $\Delta_{n+1,n+2}$ is thus bounded by a constant (less than one) times itself, and so must be zero. But then all of the other $\Delta g_{ij}$’s must also be zero, so our graphon is multipodal.
4.5. Graphons with $(n,2)$ symmetry
Finally, we show that that the optimal graphon is symmetric in the first $n$ podes and symmetric in the last two. Let $c_{1},\ldots,c_{n+2}$ be the sizes of the various podes, let $\bar{c}$ be the average size of the first $n$ podes, and let $\Delta c_{i}$ be $c_{i}-\bar{c}$ or $c_{i}-\frac{1-n\bar{c}}{2}$, depending on whether we are talking about the first $n$ podes or the last two. Let $W_{i}$ be the worth of columns in the $i$-th pode. There are five kinds of rectangles, namely $I_{i}\times I_{j}$ with $i=j\leq n$, with $i<j\leq n$ or $j<i\leq n$, with $i\leq n<j$ or $j\leq n<i$, with $i=j>n$, and finally with $\{i,j\}=\{n+1,n+2\}$. In each class, let $\bar{g}_{ij}$ be the average value of the graphon and let $\Delta g_{ij}=g_{ij}-\bar{g}_{ij}$. We also refer to $g_{n+1,n+2}$ as $p$.
A key fact is that all of the entries in the first $n$ columns are exponentially close to 0 or 1. Meanwhile, the Euler-Lagrange equations for $g_{n+1,n+2}$ say that
(107) $H^{\prime}(p)$ $\approx$ $\alpha+\beta\sum_{i=1}^{n}c_{i},$
(108) $\alpha$ $\approx$ $H^{\prime}(p)-\beta\sum_{i=1}^{n}c_{i},$
where $\approx$” means “equal up to exponentially small corrections”.
Now suppose that $i$ and $j$ are indices less than or equal to $n$. Since all of the entries $g_{ik}$ and $g_{jk}$ are exponentially close to 0 or 1, the entropy contribution to $W_{i}$ or $W_{j}$ is exponentially small. The coefficient of $\alpha$ is $\sum_{k\neq i}c_{k}=1-c_{i}$, while the coefficient of $\beta/2$ is the integral of the graphon over everything that doesn’t involve the $i$-th pode. The upshot is that
(109) $W_{i}$ $\approx$ $-\alpha(1-c_{i})-\frac{\beta}{2}(e-2c_{i}(1-c_{i}))$
(110) $=$ $-\left(\alpha+\frac{\beta e}{2}\right)+c_{i}(\alpha+\beta)-c_{i}^{2}\beta,$
with a similar result for $W_{j}$. Taking the difference gives
(111) $0$ $=$ $W_{i}-W_{j}$
(112) $\approx$ $(c_{i}-c_{j})(\alpha+\beta-\beta(c_{i}+c_{j}))$
(113) $\approx$ $(c_{i}-c_{j})(H^{\prime}(p)+\beta(c_{n+1}+c_{n+2}-c_{i}-c_{j})),$
where we have used the fact that $\sum_{i=1}^{n}c_{i}=1-c_{n+1}-c_{n+2}$. However, $c_{n+1}$ and $c_{n+2}$ are close to $\frac{1-n\bar{c}}{2}$, while $c_{i}$ and $c_{j}$ are close to $\bar{c}$, so the coefficient of $\beta$ is bounded away from zero. We conclude that $c_{i}-c_{j}$ must be exponentially small. More precisely, $c_{i}-c_{j}$ must be exponentially smaller than the largest $|\Delta g_{ik}|$ or $|\Delta g_{jk}|$. A similar argument shows that $c_{n+1}-c_{n+2}$ is also exponentially smaller than the largest $|\Delta g|$.
We now look at the Euler-Lagrange equations for $g_{i\ell}$ and $g_{j\ell}$, where $\ell$ is different from $i$ or $j$. The difference between $G(x,y)$ in $I_{i}\times I_{\ell}$ and $I_{j}\times I_{\ell}$ is
(115) $\sum_{k=1}^{n+2}c_{k}g_{\ell k}(g_{ik}-g_{jk})$ $=$ $\sum_{k=1}^{n+2}c_{k}g_{\ell k}(\Delta g_{ik}-\Delta g_{jk})$
$+(c_{i}g_{\ell i}-c_{j}g_{\ell j})(\bar{g}_{ii}-\bar{g}_{ij}).$
The first line is of the order of the largest $\Delta g$. The second line has a similar contribution from the difference of $\Delta g_{\ell i}$ and $\Delta g_{\ell j}$, plus a contribution of order $c_{i}-c_{j}$. But then
(116) $H^{\prime}(g_{i\ell})-H^{\prime}(g_{j\ell})=\beta(G_{i\ell}-G_{j\ell}),$
which is $\beta$ times a linear combination of $\Delta g$’s and $c_{i}-c_{j}$. Since $H^{\prime\prime}$ is enormous on the interval from $g_{i\ell}$ to $g_{j\ell}$ (both of which are exponentially close to 1), $\Delta g_{i\ell}-\Delta g_{j\ell}$ is bounded by a tiny combination of other $\Delta g$’s and $\Delta c$’s.
Repeating this argument for $g_{ii}-g_{jj}$ and for $g_{n+1,n+1}-g_{n+2,n+2}$, we get that
  • The biggest $\Delta c_{i}$ is bounded by a tiny constant times the biggest $\Delta g$.
  • The biggest $\Delta g$ is bounded by a tiny constant times the biggest $\Delta c$.
We conclude that all of the $\Delta c$’s and $\Delta g$’s are zero.
We have determined the form of the optimal graphon $g$ at $(e,t)$. Noting that the Boltzmann entropy $\mathbb{B}(e,t)$ equals the Shannon entropy $S(g)$ of the optimal graphon $g$, all statements about $S$ are easily converted into statements about $\mathbb{B}$.
Finally, we must show that the values of $g$ on each rectangle, and the sizes of the different podes, are analytic functions of $(e,t)$. This follows from a general principle in algebraic geometry, which in turn is essentially just the implicit function theorem. Within the product of the Razborov triangle and the finite-dimensional space of graphons with $(n,2)$ symmetry, the set of optimal graphons is a 2-dimensional analytic variety, cut out by the analytic Euler-Lagrange equations. As long as the tangent space does not degenerate, we can write all but two of the variables as analytic functions of the last two, which we can choose to be $(e,t)$.
As with Theorem 14, we have actually proven that an optimal graphon in $\mathcal{W}_{e,t}$ has a unique form up to group equivalence, implying
Corollary 18.
For all $(e,t)$ such that Theorem 17 applies, any graphon $g\in\mathcal{W}_{e,t}$ that is weakly equivalent to the multipodal entropy maximizer described in that theorem is actually group equivalent to the multipodal entropy maximizer.
4.6. Distinct phases and rank
Theorem 19.
Each of the phases above the scallops proven in Theorems 14 and 17 have unique optimal graphons with distinct symmetries and cannot be analytically continued to one another.
Proof.
The optimal graphons described by Theorem 14 have rank 2, while the optimal graphons above the $n$-th scallop described by Theorem 17 have rank $n+2$. We will construct a sequence of “order parameters,” each a polynomial in finitely many subgraph densities, to distinguish between graphons of different rank. Specifically, the $k$th order parameter is identically zero whenever the rank of the optimal graphon is k1k-1 or less, and is never zero when the rank of the graphon is $k$. Since an analytic function on a connected set that is zero on an open subset is zero everywhere, there cannot be an analytic path connecting the (k2)(k-2)-nd scallop (where the graphon has rank $k$ and the order parameter is nonzero) to the previous scallops or to the $A(2,0)$ phase, where the order parameter is zero. In other words, the phases above the different scallops are all distinct.
Newton’s identities relate the determinant of a $k\times k$ matrix $A$ to the traces of $A^{j}$ for $j=1,2,\ldots,k$. For instance, if we let $t_{j}=\mathrm{Tr}(A^{j})$, then the determinants of small matrices are given by the formulas
(117) $\det(A)=\begin{cases}(t_{1}^{2}-t_{2})/2&k=2,\cr(t_{1}^{3}-3t_{1}t_{2}+2t_{3})/6&k=3,\cr(t_{1}^{4}-6t_{1}^{2}t_{2}+8t_{1}t_{3}-3t_{4})/24&k=4.\end{cases}$
Let $p_{k}(A)$ be the polynomial in the variables $\{t_{j}\}$ that gives the determinant of a $k\times k$ matrix $A$.
The same ideas work for arbitrary diagonalizable linear operators, for which the rank equals the number of nonzero eigenvalues, counted with multiplicity. If we evaluate $p_{k}$ on any diagonalizable trace-class operator, we get zero if the rank of the operator is less than $k$ and the product of the nonzero eigenvalues (counted with multiplicity) if the rank is equal to $k$. The key algebraic fact is that, for operators of rank $k$ or less, we have
(118) $t_{j}=\lambda_{1}^{j}+\cdots+\lambda_{k}^{j},$
where some of the eigenvalues $\lambda_{i}$’s may be zero, and $p_{k}$ computes $\lambda_{1}\cdots\lambda_{k}$, which is nonzero precisely when there are $k$ nonzero eigenvalues (counted with multiplicity).
In particular, we can apply these formulas to graphons. (Graphons are always diagonalizable, being symmetric and trace class.) For instance, the expression $(t_{1}^{4}-6t_{1}^{2}t_{2}+8t_{1}t_{3}-3t_{4})/24$, where now $t_{j}=\mathrm{Tr}(g^{j})$, gives zero if the rank of the graphon $g$ is less than 4 and gives a nonzero number if the rank is equal to 4.
When $j>2$, $t_{j}$ is the density of $j$-gons. The problem is that we cannot realize $t_{1}$ and $t_{2}$ as subgraph densities, so we cannot assume a priori that $t_{1}$ and $t_{2}$ are analytic functions of $(e,t)$ in each phase. To get around this problem, we define our $k$th order parameter to be $p_{k}(g^{3})$. This is still a polynomial in $\{t_{j}\}$, only now $j$ ranges from $3$ to $3k$ in steps of 3. In particular, $t_{j}$ is the density of $j$-gons for each applicable $j$.
The $k$-th order parameter is then zero if $g^{3}$ has rank less than $k$ and is nonzero if $g^{3}$ has rank $k$. But $g^{3}$ has the same rank as $g$, so we are actually testing the rank of $g$.
In summary: the $k$th order parameter is an analytic function of $(e,t)$ in each phase, being built from subgraph densities. It is identically zero on the regions above the 0th, 1st, $\ldots$, (k3)(k-3)rd scallops but is never zero on the region above the (k2)(k-2)nd scallop. Thus the region above the (k2)(k-2)nd scallop is in a different phase from the regions above all the previous scallops. Each scallop has its own unique phase.
5. Proof of Theorem 4
Once again we prove a slightly stronger version of the theorem stated in the introduction.
Theorem 20.
For each fixed $e\in(0,1)$ and all $t$ sufficiently close to (but below) $e^{3/2}$, the optimal graphon with edge/triangle densities $(e,t)$ is unique and bipodal. Asymptotically, the Boltzmann entropy scales as $(e^{3/2}-t)\ln(e^{3/2}-t)$ and the Lagrange multiplier $\beta$ scales as $\ln(e^{3/2}-t)$.
Proof.
We follow the same overall roadmap as the proofs of Theorems 14 and 17. Specifically,
  1. (1)
    Using the proximity to the upper boundary, we break $[0,1]$ into two large podes $I_{1}$ and $I_{2}$ and a small exceptional set $I_{3}$ such that $g_{x}$ is $L^{2}$-close to the indicator function of $I_{1}$ when $x\in I_{1}$ and is $L^{2}$-close to zero when $x\in I_{2}$.
  2. (2)
    Equating the worths of $g_{x}$ when $x\in I_{1}$ to those of $g_{x}$ when $x\in I_{2}$, we determine that $\beta/\alpha\approx-2/e$. The multiplier $\beta$ is large and negative, while $\alpha$ is large and positive.
  3. (3)
    Maximizing $W(C)$ for an arbitrary $C:[0,1]\to[0,1]$, we show that every column is close to a typical column in the first or second pode. After reassigning points, $I_{3}$ is then empty. The control this gives us on $G(x,y)$ shows that $g(x,y)$ is everywhere exponentially close to 0 or 1.
  4. (4)
    Bounding the fluctuations in each rectangle by multiples of the fluctuations in other rectangles to show that all fluctuations are in fact zero. In other words, our optimal graphon is exactly bipodal with values that are exponentially close to 0 or 1.
  5. (5)
    Using Theorem 13 to eliminate the possibility that some points $(e,t)$ might have entropy maximizers without well-defined Lagrange multipliers $(\alpha,\beta)$.
Step 1 is identical to what we have done before. There is a unique reduced graphon at $(e,e^{3/2})$, namely the equivalence class of a graphon $g_{0}$ that is 1 on $I_{1}\times I_{1}$ and zero elsewhere, where $I_{1}$ is a pode of size $\sqrt{e}$. Every graphon with $t$ close to $e^{3/2}$, and in particular any entropy-maximizing graphon, must be $L^{2}$ close to $g_{0}$. This means that for all $x$’s outside of a set of small measure, $g_{x}$ is $L^{2}$-close to the corresponding column of $g_{0}$. This also implies that $G(x,y)$ is close to $\sqrt{e}$ when $x$ and $y$ are both in $I_{1}$ and is close to zero when either is in $I_{2}$.
The worth of a column that is nearly zero is of course nearly zero. The worth of a column that is nearly 1 on $I_{1}$ and nearly zero elsewhere is approximately
(119) $2\alpha\sqrt{e}-\beta e^{3/2}.$
Since all columns must have the same worth, we must have $\beta/\alpha\approx-2/e$. The Lagrange multipliers $\alpha$ and $\beta$ diverge at the same rate as we approach the boundary, with $\alpha\to\infty$ and $\beta\to$-$\infty$.
Now consider an arbitrary function $C:[0,1]\to[0,1]$. Let $a$ be the average of $C(y)$ on $I_{1}$ and let $b$ be the average on $I_{2}$. Using the approximation that $G$ is $L^{2}$-close to $\sqrt{e}$ times the indicator function of $I_{1}\times I_{1}$, we get that
(120) $W(C)\leq 2\sqrt{e}H(a)+2(1-\sqrt{e})H(b)-2\alpha\sqrt{e}a-2\alpha(1-\sqrt{e})b-\beta e^{3/2}a^{2},$
with equality if $C$ is constant on $I_{1}$ and constant on $I_{2}$. Since $\alpha$ is large and positive, we must have $b$ exponentially close to 0. Setting $b\approx 0$, our worth is then approximately
(121) $\alpha\sqrt{e}(a^{2}-a).$
This is of course maximized at the endpoints $a=1$ and $a=0$, being negative when $a\in(0,1)$. In other words, any worth-maximizing column must either have $a\approx 1$ and $b\approx 0$, and so must be close to the columns in $I_{1}$, or $a\approx 0$ and $b\approx 0$, and so must be close to the columns in $I_{2}$. Reassigning the points of $I_{3}$ to $I_{1}$ or $I_{2}$ accordingly, we obtain a situation where $I_{3}$ is empty.
To constrain the fluctuations in $g(x,y)$ in each rectangle, we recall the variational equations
(122) $H^{\prime}(g(x,y))=\alpha+\beta G(x,y).$
Since $G(x,y)\approx 0$ or $\sqrt{e}$, depending on which quadrant we are in, this implies that $g(x,y)$ is exponentially close to 1 on $I_{1}\times I_{1}$ and exponentially close to 0 on $I_{1}\times I_{2}$ and $I_{2}\times I_{2}$. In particular, $H^{\prime\prime}(g)$, which scales as the larger of $1/g$ and 1/(1g)1/(1-g), is much larger than $|\beta|$. Looking at the change in the left hand side and right hand side of this equation within a single quadrant, we see that $H^{\prime\prime}(g)$ times the maximum fluctuation within any quadrant is of the same order as $|\beta|$ times the maximum fluctuation within any quadrant. But that means that the maximum fluctuation is bounded by a small multiple of itself, and so must be zero. Our graphon is exactly bipodal.
Finally, we do some calculations in the space of bipodal graphons. Let $g_{11}$, $g_{12}$ and $g_{22}$ be the values of the optimal graphon on $I_{1}\times I_{1}$, $I_{1}\times I_{2}$ and $I_{2}\times I_{2}$, respectively. We treat $g_{11}$, $g_{12}$ and $g_{22}$ as free variables and adjust the width of $I_{1}$ to keep the edge density fixed. To leading order, $e^{3/2}-t$ is a linear function of $(1-g_{11},g_{12},g_{22})$. However, $1-g_{11}$, $g_{12}$ and $g_{22}$ all scale as exponents of $\beta$, so $\beta$ must scale as $\ln(e^{3/2}-t)$. The entropy goes as $(1-g_{11})\ln(1-g_{11})-g_{12}\ln(g_{12})-g_{22}\ln(g_{22})$, which then scales as $(e^{3/2}-t))\ln(e^{3/2}-t)$.
The analyticity of $g$ as a function of $(e,t)$ follows from the same argument as in the proof of Theorem 17, only with the space of $(n,2)$ symmetric graphons replaced by the space of bipodal graphons.
Finally, as with Theorems 14 and 17, we note that we have determined the entropy maximizer up to group equivalence, implying
Corollary 21.
Let $(e,t)$ be such that Theorem 20 applies. Then any graphon $g\in\mathcal{W}_{e,t}$ that is weakly equivalent to the bipodal maximizer defined in the theorem is actually group equivalent to that bipodal maximizer.
6. ERGM-invisibility
Using Lagrange multipliers is superficially similar to studying an exponential random graph model (ERGM; see Section 6 in [11], or [15], or [13] for a relevant introduction), where one considers an ensemble of all graphs on $n$ vertices, where the probability of a given graph $G$ is proportional to
(123) $\exp\left(-n^{2}\left[\alpha\varepsilon(G)+\frac{\beta}{3}\tau(G)\right]\right).$
Here $\varepsilon(G)$ and $\tau(G)$ are the edge and triangle densities of the graph $G$ and $\alpha$ and $\beta$ are variables which can move the distribution, expected to be narrowly peaked for large $n$. (In the literature, ERGMs are usually described in terms of parameters $\beta_{1}=-\alpha$ and $\beta_{2}=-\beta/3$, but that linear change of variables does not matter.)
ERGMs are widely used to model real-world networks, with $n$ necessarily finite. See [20], [21] and their bibliographies for references relevant to the current discussion. By 2010 there was literature noting that fitting of parameters to data had various problems, and in [13] the recent LDP for $\mathbb{G}(n,p)$ graphs [14] was applied to see if that would help. See [9] and the introduction in [13] for background. [13] illuminated some issues but left some others unresolved. Our treatment of the Boltzmann entropy can help, as follows.
The $n\to\infty$ limit of an ERGM with parameters $(\alpha,\beta)$ can be understood in terms of graphons and the function
(124) $\Psi(\alpha,\beta)=\max_{g}\left(S(g)-\alpha\varepsilon(g)-\frac{\beta}{3}\tau(g)\right)\equiv\max_{g}F(g).$
The LDP relates the graphon that maximizes $F(g)$ on the right hand side of (124), for the given values of $\alpha$ and $\beta$, to typical large graphs in the ensemble. Note that $\Psi(\alpha,\beta)$ is the Legendre transform of the Boltzmann entropy $\mathbb{B}(e,t)$:
(125) $\Psi(\alpha,\beta)=\max_{e,t}\left(\mathbb{B}(e,t)-\alpha e-\frac{\beta}{3}t\right).$
If the Boltzmann entropy function were convex and the Razborov triangle were convex, then the Legendre transform (125) would be invertible within each phase [52]. If that were true, we could tune $\alpha$ and $\beta$ within each phase to get whatever values of $(e,t)$ we wanted. In statistical mechanics, this ability to switch back and forth between fundamental variables and conjugate variables is called equivalence of ensembles.
With graphs, the Boltzmann entropy function is not convex and neither is the Razborov triangle. There is no equivalence of ensembles; the Legendre transform (125) is not invertible. Specifically, there are many values of $(e,t)$ for which there do not exist any values of $(\alpha,\beta)$ whose $F$-maximizing graphons have edge/triangle densities $(e,t)$. We call such points $(e,t)$ ERGM-invisible. We can still understand $\Psi(\alpha,\beta)$ and the phases of an ERGM by studying $\mathbb{B}(e,t)$, since $\Psi$ is still the Legendre transform of $\mathbb{B}$. However, we cannot understand $\mathbb{B}(e,t)$, or graphs with general densities $(e,t)$, by studying $\Psi(\alpha,\beta)$.
It has long been known that all points with $t$ greater than $e^{3}$, and even moderately less than $e^{3}$, are ERGM-invisible [13]. The only points off the Erdős-Rényi curve that might be ERGM-visible lie close to the lower boundary of the triangle (Figure 1). We now show that, because of the nonconvexity of the Razborov triangle, most of those are also ERGM-invisible.
Theorem 22.
If $n$ is a positive integer and $\frac{n}{n+1}<e<\frac{n+1}{n+2}$, and if $t$ is sufficiently close to the minimum triangle density $t_{0}$, then $(e,t)$ is ERGM-invisible.
Proof.
Fix a value of $e$ strictly between $\frac{n}{n+1}$ and $\frac{n+1}{n+2}$. The points with edge density $e$ and triangle density just above the minimum must have large values of the Lagrange multipliers $\alpha$ and $\beta$. However, the scallop itself is concave down, so for positive values of $\beta$, the linear function $\alpha e+\beta t$ is greater at one or both of the neighboring cusps (at edge density $\frac{n}{n+1}$ and $\frac{n+1}{n+2}$) than near the interior of the scallop. For large positive values of $\beta$ (and correspondingly large negative values of $\alpha$), this difference is greater than the bounded difference in Shannon entropy between the graphons described by Theorem 17 and the zero-entropy graphons at the cusps. Since large values of $\beta$ correspond to small values of $\Delta t$, we conclude that, for all sufficiently small values of $\Delta t$, the point $(e,t_{0}+\Delta t)$ is ERGM-invisible. ∎
A similar result applies at the top of the Razborov triangle.
Theorem 23.
For each $e\in(0,1)$ and for all $t$ sufficiently close to (and less than) $e^{3/2}$, $(e,t)$ is ERGM-invisible.
Proof.
Near the top boundary, $\alpha$ is large and positive while $\beta$ is large and negative. However, the boundary curve $t=e^{3/2}$ is concave up, so we can decrease $\alpha e+\beta t/3$ by moving to one endpoint $(0,0)$ or the other $(1,1)$. Whenever $\alpha$ and $\beta$ are big enough in magnitude, in other words whenever we are close enough to the top boundary, these gains swamp any changes in $S(g)$. Either the constant graphon $g=0$ or the constant graphon $g=1$ yields a larger value of $F$ than the optimal graphon at $(e,t)$. ∎
The statement of Theorem 23 is not new; see [13, Theorem 6.2]. However, the simplicity of the proof gets to the heart of why points in this region are ERGM-invisible.
Considering the Razborov triangle as a whole, only a few pieces are ERGM-visible. The Erdős-Rényi curve $t=e^{3}$ is ERGM-visible. Since $d\varepsilon$ and $d\tau$ are collinear at constant graphons, each point on the Erdős-Rényi curve actually corresponds to an infinite set of $(\alpha,\beta)$ values. A neighborhood of each cusp is ERGM-visible; that’s what you get when $\beta$ is large and positive. As far as we can tell, a neighborhood of the flat part of the lower boundary is also ERGM-visible. In short, ERGMs can see homogeneous Erdős-Rényi or $A(n,0)$ structures, where every vertex looks like every other vertex. But that’s all.
In this paper, we have shown how a homogeneous constraint on the total number of edges and triangles leads to inhomogeneous $F(1,1)$ or $C(n,2)$ structures, where vertices in one set of podes look very different from vertices in another set of podes. We previously showed similar behavior near the Erdős-Rényi curve. However, ERGMs cannot see this spontaneous emergence of inhomogeneity, as the portions of the Razborov triangle where such emergent inhomogeneity occurs are all ERGM-invisible.
This is not to deny the real successes that ERGMs have had. Starting with [14] and [13], definitions were given of phases in the $(\alpha,\beta)$ plane. The existence of phases and phase transitions was proven quickly, far faster than for the analogous problems discussed in this paper. Those are important triumphs. Fundamentally, the two approaches to studying large graphs represent different parts of applied mathematics, with different goals.
Acknowledgment. The work of L.S. was partially supported by the National Science Foundation under grant DMS-2113468.